### v2

Abstract
Deep learning, a branch of artificial intelligence, is a computational model that uses multiple layers of interconnected units (neurons) to learn intricate patterns and representations directly from raw input data. Empowered by this learning capability, it has become a powerful tool for solving complex problems and is the core driver of many groundbreaking technologies and innovations. Building a deep learning model is a challenging task due to the algorithm’s complexity and the dynamic nature of real-world problems. Several studies have reviewed deep learning concepts and applications. However, the studies mostly focused on the types of deep learning models and convolutional neural network architectures, offering limited coverage of the state-of-the-art of deep learning models and their applications in solving complex problems across different domains. Therefore, motivated by the limitations, this study aims to comprehensively review the state-of-the-art deep learning models in computer vision, natural language processing, time series analysis and pervasive computing. We highlight the key features of the models and their effectiveness in solving the problems within each domain. Furthermore, this study presents the fundamentals of deep learning, various deep learning model types and prominent convolutional neural network architectures. Finally, challenges and future directions in deep learning research are discussed to offer a broader perspective for future researchers.
keywords:
[…]
1 Introduction
Deep learning has revolutionized many applications across a variety of industries and research. The application of deep learning can be found in healthcare Shamshirband et al. (2021), smart manufacturing Wang et al. (2018b), robotics Pierson and Gashler (2017), cybersecurity Dixit and Silakari (2021) etc., solving challenging and complex problems such as disease diagnosis, anomaly detection, object detection and malware attack detection. Deep learning is a subset of machine learning that learns from data using artificial neural networks. An artificial neural network is a computational model that imitates the working principles of a human brain. The computational models are composed of an input layer which receives the input data, multiple processing layers that learn the representation of data and the output layer which produces the output of the model.
Prior to the reintroduction of deep learning (DL) into the research trend, pattern recognition tasks involved a transformation of the raw input data such as pixel values of an image into a feature vector that represents the internal representation of the data. The feature vector can be used by a machine learning model to detect or classify patterns in the data. This process requires feature engineering and considerable domain knowledge to design a suitable feature representation. With deep learning, this cumbersome process can be performed automatically whereby at each processing layer known as hidden layers, the internal representation of the input data is learned or extracted in a hierarchical manner. The first layer learns the presence of basic primitive features such as edges, dots, lines etc. The second layer learns patterns or motifs by recognizing the combinations of the edges, dots and lines, and the subsequent layers combine the motifs to produce more sophisticated features that correspond to the input data. This feature learning process takes place in the sequence of hidden layers until the prediction is finally produced.
Several studies have been conducted to discuss the concept and application of deep learning in the last few years, as listed in Table 1. The studies addressed or focused on several aspects of deep learning, such as types of deep learning models, learning approaches and strategies, convolutional neural network (CNN) architectures, deep learning applications and challenges. In Dong et al. (2021), the authors provided fundamentals of deep learning and highlighted different types of deep learning models, such as convolutional neural networks, autoencoder and generative adversarial networks. Then, the applications of deep learning in various domains are discussed, and some challenges associated with deep learning applications are presented. Another survey Talaei Khoei et al. (2023) provided a comprehensive analysis of supervised, unsupervised and reinforcement learning approaches and compared the different learning strategies such as online, federated and transfer learning. Finally, the current challenges of deep learning and future direction are discussed.
In Alzubaidi et al. (2021), the authors provided a comprehensive review of the popular CNN architectures used in computer vision tasks, highlighting their key features and advantages. Then, the applications of deep learning in medical imaging and the challenges are discussed. A similar survey is reported in Alom et al. (2019), where the different supervised and unsupervised deep learning models are highlighted, and the popular CNN architectures are compared and discussed. In another survey Pouyanfar et al. (2018), the authors focused on the applications of deep learning in computer vision, natural language processing and speech and audio processing. The different types of deep learning models are also discussed. In Sarker (2021), the authors focused on the different types of deep learning models and provided a summary of deep learning applications in various domains.
Despite the existing surveys on deep learning that offer valuable insights, the increasing amount of deep learning applications and the existing limitations in the current studies motivated us to explore this topic in depth. In general, to the best of our knowledge, no survey paper focuses on the emerging trends in state-of-the-art applications and the current challenges associated with deep learning. Furthermore, the surveys do not discuss the issues and how deep learning addresses them by highlighting the key features and components in the models. Also, most surveys either ignore or provide minimal coverage of the fundamentals of deep learning, which is crucial for understanding the state-of-the-art models. The main objective of this paper is to present the most important aspects of deep learning, making it accessible to a wide audience and facilitating researchers and practitioners in advancing and leveraging its capabilities to solve complex problems across diverse domains. Specifically, we present the fundamentals of deep learning and the various types of deep learning models, including popular deep learning architectures. Then, we discuss the progress of deep learning in state-of-the-art applications, highlighting the key features of the models and their problem-solving approaches. Finally, we discuss the challenges faced by deep learning and the future research directions.
Table 1: Summary of related works.
Reference Focus Concepts not covered
Dong et al. (2021) A short review of the fundamentals of DL networks and discusses different types of neural networks, DL applications and challenges. Lack of analysis of CNN architectures and limited coverage of deep learning fundamentals.
Talaei Khoei et al. (2023) Discusses the learning approaches (supervised, unsupervised and reinforcement learnings), learning strategies, and DL challenges Lack of fundamentals of deep learning, CNN architectures and DL applications.
Alzubaidi et al. (2021) Discusses different types of DL networks, CNN fundamentals and architectures, DL challenges and medical imaging applications Limited discussion on DL applications such as natural language processing and time series analysis.
Alom et al. (2019) A short review of the fundamentals of neural networks and discusses different types of DL networks, CNN architectures and applications. Limited discussion on DL applications and no discussion of DL challenges.
Pouyanfar et al. (2018) Discusses different types of DL networks and DL applications and challenges Lack of analysis of CNN architectures and limited coverage of deep learning fundamentals.
Sarker (2021) Discusses different types of DL networks and provides a summary of DL applications Lack of fundamentals of deep learning, analysis of CNN architectures and limited discussion on DL applications.
The remainder of this paper is organized as follows: Section 2 describes the fundamentals of deep learning which includes layers and attention mechanisms, activation functions, model optimization and loss functions, and regularization methods. Section 3 presents the types of deep learning models, including the CNN architectures. Section 4 discusses the state-of-the-art applications of deep learning. Section 5 discusses the challenges and future directions in the field of deep learning. The conclusion is given in Section 6.
2 Fundamentals of Deep Learning
This section describes the fundamental concepts such as layer types, activation functions, training algorithms and regularization methods to provide a comprehensive understanding of the underlying principles in advancing the field of deep learning.
2.1 Layers
A deep learning model is characterized by having numerous hidden layers. The hidden layers are responsible for learning and extracting complex features from the input data. A hidden layer is composed of an arbitrary number of neurons which serves as the fundamental building block of a neural network as shown in Fig. 1. A neuron consists of an arbitrary number of inputs, each associated with a weight, which controls the flow of information into the neuron during the forward pass. The flow of information, or forward pass, involves the computation of summation of the weighted input, followed by the application of a transformation function to the weighted sum. Consider a neuron $z_{i}$ at layer $l$, receives an input vector $a_{j}^{(}l-1)$, the computation of the neuron is defined as
$z_{i}^{l}=\sum_{j=0}^{d}w_{i,j}^{l-1}\cdot a_{j}^{l-1}$
$a_{i}^{l}=g(z_{i}^{l})$
where $w_{ij}$ is the set of weights connecting the inputs to the neuron and g is the transformation function also known as activation function. A hidden layer wherein each neuron is connected to all neurons of the previous layer is known as fully-connected layer.
Figure 1: A graphical representation of a neuron.
The key aspect of deep learning lies in its ability to automatically extract hierarchical features from the input data. This automatic feature extraction is performed by two specialized layers called convolutional and pooling layers. In a convolutional layer, each neuron is connected only to a local region of the input data, and the weights are shared across the input data. The weight-sharing not only significantly reduces the number of parameters of the neural network, but also allows the network to learn the same features across different spatial locations in the input LeCun et al. (2015). Fig. 2 illustrates a convolutional layer applies $3\times 3$ filter on two-dimensional image consisting of $9\times 9$ pixels. The layer convolves the input data by moving the filter across the whole input pixels, producing a set of output values called feature map. The computation of a convolutional layer $l$ is defined as
$z_{i,j,d}^{l}=\sum_{m=0}^{k_{1}}\sum_{n=0}^{k_{2}}w_{m,n,d}^{l-1}\cdot a_{i+m,j+n,d}^{l-1}$
$a_{i,j,d}^{l}=g(z_{i,j,d}^{l})$
where $w_{m,n,d}$ is the weight of $k_{1}\times k_{2}$ filter and $a_{i,j,d}^{l-1}$ is the input pixel.
Figure 2: A neuron is connected to a local region of the input data.
Pooling layers are commonly applied after successive convolutional layers to progressively reduce the spatial dimensions of the feature maps. The spatial reduction is performed by computing the summary of the local regions in the feature maps. Two common pooling operations are computing the maximum and the average of the local regions. This reduces the number of parameters while providing translation-invariant features LeCun et al. (2015).
2.2 Attention Mechanisms
One of the important concepts in pattern recognition is the ability to attend and neglect certain parts of the input data based on their importance. This is because not all parts of the input hold equal importance for making the prediction. Certain features exhibit a stronger correlation with the output while others are less relevant. In convolutional layers, all extracted features are treated uniformly, without consideration of the varying degree of the importance of the different parts of the input data. This limitation is addressed by the introduction of attention mechanism, which can dynamically assign varying levels of significance (weights) to the different features. This flexibility enables the deep learning models to prioritize the more relevant aspects of the input data, enhancing its ability to capture the intricate dependencies for accurate prediction. Given an input data $x$, the process of attending to the important components of the input is given as
$A=f(g(x),x)$
where $g$ is a composite function that performs a sequence of operations to generate the attention or the weights and f applies the generated attention $g(x)$ on the input $x$, $f=g(x)x$.
For instance, the squeeze-and-excitation (SE) attention generates the attention through five consecutive operations Hu et al. (2019). First, the input is vectorized using global average pooling. Then the vector is passed to two fully-connected layers, where the first one with ReLU activation and the second one with sigmoid activation. SE attention was a pioneer in channel attention. The attention module assigns varying weights to the channels of the feature maps. SE attention suffers from computational cost and the use of global average which may cause information loss at the spatial level. Several efforts have been made to improve SE attention. GSoP attention performs $1\times 1$ convolution on the feature maps to reduce the number of channels, and then computes the pairwise channel correlation which is used to generate the weights Gao et al. (2018). ECA attention replaced the fully-connected layers with 1d convolution to reduce the number of parameters and the computational cost Wang et al. (2020a).
Temporal attention is an attention module that focuses on specific time steps in a sequence of data such time series and video (sequence of images). In video processing such as recognizing human actions, temporal attention is used to focus on key frames at different point in time that contains crucial information for predicting the ongoing activity. Temporal adaptive module (TAM) is a temporal attention that can focus on short-term (local) information and global context information of the data Liu et al. (2021b). The composite function consists of local branch for generating attention weights and global branch for generating channel-wise adaptive kernel. First, the input feature map is squeezed using global average pooling to reduce the computational cost. Subsequently, the local branch executes two 1D convolution operations, with the first convolution using ReLU, and sigmoid activation for the second to generate the local weights. The local weights are then multiplied with the featuremap. Meanwhile, the global branch is composed of two fully-connected layers, with the first layer using ReLU and second layer employing softmax function to generate the adaptive kernel (weights). Self-attention is a form of temporal attention, initially proposed for machine translation to enable the deep learning models to attend different words in a sequence relative to other words Vaswani et al. (2023). The attention module has become a fundamental building block in various natural language processing applications. To generate the attention weights, the input (word embeddings) is transformed by linear projection to compute query, key and value. Then, dot product between query and key is computed, and the resultant is normalized by the square root of the size of the key. Finally, the attention weights are obtained by applying the softmax function. Self-attention is the fundamental building block of the Transformer architecture, a key deep learning model in natural language processing.
Spatial attention focuses on specific regions or spatial location of the input data, enabling the deep learning models to selectively emphasize and ignore certain features. In the context of computer vision, spatial attention is crucial in capturing the spatial relationships and context within an image for accurate prediction. Attention gate is a spatial attention that can identify and focus the salient regions and suppress feature responses of the insignificant ones. The composite function consists of ReLU activation followed by $1\times 1$ convolution to reduce channel dimension of the feature maps to a singular feature map. Finally, sigmoid is applied to the feature map to generate attention weights Oktay et al. (2018). The self-attention in the standard Transformer is not effective in handling image data due to its inherent sequential processing nature and lacks the ability to capture spatial dependencies and local patterns. To address this limitation, the Vision Transformer (ViT) treats images as a sequence of non-overlapping patches. A similar computational pipeline is used to generate the attention weights, the sequence of patches is transformed by linear projection to compute the query, key and value Dosovitskiy et al. (2021). The same operations are employed to generate the attention weights. Self-attention is computationally costly due to its quadratic complexity especially dealing with image data. To reduce the complexity, two learnable linear layers, independent of the input data are adopted as the key and value vectors Guo et al. (2023).
2.3 Activation Functions
The role of the activation function is to transform the weighted sum into a more classifiable form. This is crucial to the learning behaviour of the deep learning model, generating non-linear relationships between the input and the output of the model. The activation function, combined with many hidden layers allow the neural network to approximate highly complex, non-linear functions. Many activation functions are available for use in neural networks, and some of the functions are shown in Fig. 3 - Fig 5. The figures show the plot of the three popular activation functions. The sigmoid is a classic example of activation function which is used in logistic regression. Sigmoid activation function maps the weighted sum to a value in the range of 0 and 1 which can be used for classification. Hyperbolic tangent is another popular choice of bounded activation function which produces an output between -1 and 1. Hyperbolic tangent has a stronger gradient, hence the neural network training often converges faster than sigmoid LeCun et al. (2012). For many years, sigmoid and hyperbolic tangent are the commonly used activation functions. Nevertheless, the activation functions suffer from vanishing gradient problem, hindering the efficient training of deep neural networks with many layers Hochreiter et al. (2001).
It was shown that neural networks with unbounded activation functions have the universal approximation property and reduce the vanishing gradient problem. Over the past years, numerous unbounded activation functions were proposed for neural networks, with softplus Dugas et al. (2000) and rectified linear unit (ReLU) Glorot et al. (2011) activation function being the notable examples. These activation functions especially ReLU has pivotal role in improving the training and performance of deep learning models. ReLU has been a cornerstone of deep learning models due to its computational efficiency and effectiveness in addressing the vanishing gradient problem. Since then, several variants of ReLU were proposed including Leaky ReLU Maas et al. (2013), sigmoid linear unit Misra (2020) and exponential linear unit Clevert et al. (2016), each offering unique advantages for building deep learning-based applications.
Figure 3: Sigmoid activation function.
Figure 4: Hyperbolic tangent activation function.
Figure 5: Rectified linear unit activation function.
2.4 Parameter Learning and Loss Functions
The weights (parameters) of deep learning models are optimized using an optimization algorithm called gradient descent. However, it has to be noted that gradient descent is a generic algorithm which can be used to solve a wide range of optimization problems. In general, gradient descent finds the optimal weights by iteratively updating the weights such that the weights will result in a minimum prediction error over all instances in the training set. The prediction error is quantified by a loss function. For classification problems, the commonly used loss function is the negative log-likelihood loss or cross entropy loss, while the square loss and absolute loss are used for regression problems Wang et al. (2022). The weight update is defined as
$w_{i,j}^{l}=w_{i,j}^{l}-\alpha\nabla_{w_{i,j}}$
where $\alpha$ is a hyperparameter called learning rate and $\nabla_{w_{i,j}}$ is the gradient or the derivative of the loss function $J$ with respect to the weight $\frac{\partial J}{\partial w_{i,j}^{l}}$.
The gradient can be computed across all instances of the training set, an approach known as batch gradient descent. However, this approach may not guarantee a convergence to the optimal solution as the gradient is the same for every weight update. An alternative approach is to perform the weight update on the basis of a single instance, but the approach results in a noisy gradient and becomes computationally intensive due to the frequent weight update. A more commonly used approach is to perform the weight update over a set of training instances, known as mini-batch gradient descent. This approach strikes a balance, providing a less noisy gradient and a more stable training process.
Several efforts have been made to improve the efficiency of gradient descent. One of the earlier efforts is the inclusion of past rate of change in the weight update to speed up the training of deep learning models, the algorithm is called gradient descent with momentum Qian (1999). Another effort is to improve the training convergence by adapting the learning rate based on the occurrence of the features Duchi et al. (2011). A more recent work utilizes both adaptive learning rate and momentum to improve the training efficiency and convergence of deep learning models Kingma and Ba (2017).
2.5 Regularization Methods
Regularization methods are employed to prevent overfitting in deep learning models and improve their generalization performance. Early stopping is a method that can detect the onset of overfitting during training by continuously monitoring the validation error. The model is considered overfitting if the validation error starts to increase at some point of the training while the training error is decreasing. However, detecting the onset of overfitting during the training of deep learning models is challenging due to the inherent stochasticity and the presence of noisy data. Several stopping criteria can be considered such as using a threshold to check if the decrease of (average) validation error is significant and count the number of successive increases of validation error Prechelt (2012).
Dropout is a regularization method that randomly switching off some neurons in the hidden layers during training with a predefined drop probability (dropout rate) Srivastava et al. (2014). Dropout has the effect of training and evaluating exponentially many different deep learning models. The dropout rate is a hyperparameter that needs to be carefully tuned to balance regularization and model capacity. Different ranges of dropout rate have been suggested. The original author suggested a dropout rate between 0.5 and 0.8 Srivastava et al. (2014) while others recommended a lower dropout rate between 0.1 and 0.3 Park and Kwak (2017). Also, it has been suggested a low dropout rate due to the exponential increase in the volume of training data Liu et al. (2023b).
Parameter norm penalty is a regularization method that adds a penalty term consisting of the network’s weights to the loss function. During the training, the penalty term discourages large weight values and hence, constraining the model’s capacity and reducing the chance of overfitting. The common penalty terms are $L^{1}$ norm penalty Tibshirani (1996), $L^{2}$ norm penalty, also known as weight decay and a combination of $L^{1}$ and $L^{2}$ Zou and Hastie (2005). An adaptive weight decay is proposed allowing the regularization strength for each weight to be dynamically adjusted Nakamura and Hong (2019).
Despite the advantages of the mini-batch gradient descent, each mini-batch may comprise data from different distributions. Furthermore, the data distribution may change after each weight update, which could slow down the training process. Batch normalization overcomes this issue by normalizing the summed input to a neuron over a mini batch of training instances Ioffe and Szegedy (2015). An alternative method is to perform normalization across the neurons instead of the mini batch, a method known as layer normalization Ba et al. (2016). Layer normalization is applicable in recurrent neural network and overcomes the dependencies on the mini batch size.
3 Types of Deep Learning
Deep learning models can be categorized into deep supervised learning and deep unsupervised learning.
3.1 Deep Supervised Learning
Deep supervised models are trained with a labelled dataset. The learning process of these models involve calculating the prediction error through a loss function and utilizing the error to adjust the weights iteratively until the prediction error is minimized. Among the deep supervised models, three important models are identified namely multilayer perceptron, convolutional neural network and recurrent neural network.
Multilayer perceptron is a neural network model with one or more hidden fully-connected layers stacked between the input and output layers as shown in Fig. 6. The width (number of neurons) of the hidden layers and the depth (number of layers) of the network influence the model’s ability to learn patterns in the data. Specifically, the width determines the model’s ability to learn complex features while the depth allows the model to learn hierarchical representations of the data. Nevertheless, studies showed that a multilayer perceptron with a single hidden layer can approximate any continuous function Cybenko (1989), Hornik et al. (1989). Multilayer perceptron is effective in various industries and applications from healthcare to finance Widrow et al. (1994). However, multilayer perceptron requires the input data to be structured in a one-dimensional format e.g. tabular data, making it less suitable for unstructured data such as image, text and speech. To leverage multilayer perceptron for unstructured data, a feature extraction or transformation into structured data is necessary.
Figure 6: A fully-connected neural network.
Recurrent Neural Network is a neural network model that leverages the sequential information and memory through the use of recurrent connections, allowing it to effectively process data such as time series, text, speech and other sequential patterns. As shown in Fig. 7, a recurrent neural network is characterized by the recurrent connection which enables the network to loop back and use internal state from previous time step to the next time step. The internal state is parameterized by a set of weights which is shared across the sequence of the data. The training of recurrent neural networks suffers from the issue of vanishing gradient due to the challenges of propagation of gradients over a long sequence of data. Variants of recurrent neural networks are introduced to overcome the problem of vanishing gradient such as long short-term memory Hochreiter and Schmidhuber (1997) and gated recurrent memory Cho et al. (2014). The improved recurrent neural networks introduce memory cell and gating mechanisms to retain and discard information in every time step, allowing for more effective learning dependencies in long sequence. The network architecture can be built using fully-connected and convolutional layers Shi et al. (2015).
Figure 7: A neural network with recurrent connection.
Convolutional Neural Network (CNN) is a neural network model that preserves and leverages the spatial local information in the data through the use of convolutional layers. Fig. 8 shows a typical architecture of convolutional neural network which comprises of convolutional, pooling and fully-connected layers. The convolutional and pooling layers are stacked alternately to automatically extract salient features in a hierarchical manner. The extracted features are then fed to fully-connected layers to predict the outputs. The final feature maps need to be converted to one-dimensional vector before they are fed to the fully-connected layers. The conversion can be performed by flattening the feature maps. CNN architecture is crucial in increasing the performance of the prediction, as it is designed to efficiently extract the feature representation of the input data, enabling more accurate and robust pattern recognition. Over the last decade, several CNN architectures have been proposed, whereby the focus of the improvements has been on enhancing the feature learning capabilities and addressing challenges such as vanishing gradient and diminishing feature reuse.
Figure 8: A neural network with convolutional and pooling layers followed by fully-connected layers.
AlexNet is among the first CNN models that gained widespread recognition and success, marking a significant achievement in the field of deep learning for computer vision tasks Krizhevsky et al. (2012). The model consists of five convolutional layers with max-pooling operation performed after the first and second convolutional layers, followed by three fully-connected layers. The first and second convolutional layers utilize a filter size of 11×11 and 5×5 respectively, and $3\times 3$ filter size is used for the remaining convolutional layers. ReLU activation function is used to mitigate the vanishing gradient.
VGG-16 attempts to improve the CNN architecture by adding more convolutional layers, specifically up to 19 layers to capture more intricate feature representation from input data, followed by three fully-connected layers Simonyan and Zisserman (2015). ReLU activation function is used to reduce vanishing gradient. Unlike AlexNet, all convolutional layers utilize a small fix filter size of $3\times 3$ and max-pooling layer is added after a stack of two or three convolutional layers. This configuration allows the model to extract more discriminative features and decreases the number of parameters.
ZFNet is a classic CNN model which has a similar architectural principle as AlexNet, featuring five convolutional layers with max-pooling layers after the first and second convolution, followed by three fully-connected layers Zeiler and Fergus (2013). The significant differences are the use of smaller filter size and stride in the convolutional layers and contrast normalization of the feature maps which allows the model to capture better features and improving the overall performance.
Network-in-network introduces two innovative concepts to enhance the performance of the model Lin et al. (2014). The first was introducing a block of convolutional layers consisting of k×k convolution followed by two $1\times 1$ convolution operations. The pointwise convolutions are similar to applying multilayer perceptron on the feature maps, allowing the model to approximate more abstract feature representation. In the preceding models, the final feature maps are vectorized by flattening operation for classification by the fully-connected layers. Instead of flattening, network-in-network model calculates the spatial average of each feature map, and the resulting vector is fed to softmax function for classification. This approach is parameter-less, significantly reducing the number of parameters.
GoogleNet leverages the fact that visual data can be represented at different scales by incorporating a module which consists of multiple convolutional pipelines with different filter sizes Szegedy et al. (2014). The module known as inception utilizes three kernel sizes ($5\times 5$, $3\times 3$, $1\times 1$) to capture spatial and channel information at different scales of resolution as shown in Fig. 9. This configuration enables a more effective feature extraction at both fine-grained and coarse-grained information from input data. The model architecture utilizes the inception module at the higher layers while the traditional convolution and max-pooling block is used to extract primitive and basic features. The inception modules are stacked upon each other with maximum pooling operation is performed occasionally to reduce the spatial resolution of the feature maps. GoogleNet utilizes global average pooling to vectorize the final feature maps before passing it to a fully-connected layer for classification.
Figure 9: The inception module.
Increasing the number of layers enhances the model performance, mainly for solving complex tasks. However, training a very deep neural network is challenging due to the vanishing gradient problem, where the gradients that are used to update the network become insignificant or extremely small as they are backpropagated from the output layer to the earlier layers. A model called Highway Network overcomes this issue by introducing a gating mechanism that regulates the information flow of the layers, enabling the flow of information from the earlier layers to the later layers Srivastava et al. (2015). Consequently, this not only mitigates the vanishing gradient problem, but also renders the gradient-based training more tractable, enabling the training of very deep neural networks consisting as many as 100 layers.
The gating mechanism of Highway Network increases the number of parameters for regulating the information flow. ResNet is a CNN architecture that incorporates residual (skip) connection that allows information to bypass certain layers, mitigating the vanishing gradient problem He et al. (2015). ResNet architecture stacks residual blocks, which consists of two or three of convolutional layers with batch normalization and ReLU, and a skip connection which adds the input to the output of the final convolutional layer as shown in Fig. 10. If the input dimension does not match with the residual output dimension, a linear projection is performed by the residual connection to match the dimensions. In comparison to the gating mechanism of Highway Network, the residual connection is parameter-free, and thus does not incur additional computational costs. Furthermore, the connections are never closed whereby all information is always passed through the layers. This innovative concept enables the training of very deep neural networks boasting as many as 152 layers.
Figure 10: A residual connection.
DenseNet is another CNN architecture that overcomes the vanishing gradient problem. DenseNet follows the same approach as ResNet and Highway Network, utilizing skip connection to allow information flow from the earlier layers to later layers. However, DenseNet takes this concept one step further, by introducing a dense block consisting of multiple convolution functions (layers) with each convolution function performs batch normalization followed by ReLU and $3\times 3$ convolution. Each convolutional layer in the dense block receives feature maps from all its preceding layers, hence the connection is referred to as dense connection Huang et al. (2017). This configuration as shown in Fig. 11 maximizes information flow and preserves the feed-forward nature of the network. Dense blocks can become computationally expensive due to the increasing number of feature maps. To reduce the computational costs, a block of $1\times 1$ convolutional with batch normalization and max-pooling layers known as transition block is used to reduce the spatial dimension of the feature maps. The model architecture integrates these dense and transition blocks, stacking them alternately. The network depth can reach up to 264 layers.
Figure 11: A dense block.
Although skip connections in ResNet effectively mitigate the vanishing gradient problem, a new challenge arises in the form of diminishing feature reuse as the network becomes deeper. Diminishing feature reuse refers to the diminishing effectiveness of the previously learned feature maps in subsequent layers, impacting the final prediction. WideResNet is a CNN architecture that is based on ResNet with the aim to mitigate diminishing feature reuse problem. Instead of making the network deeper, WideResNet makes the network wider by increasing the number of channels by k factor Zagoruyko and Komodakis (2017). The increased width allows the model to capture a more diverse features, enhancing its ability to learn complex relationships in the input data.
ResNext addresses the diminishing feature reuse by capturing more efficient and diverse features of the input data. ResNext introduces a concept of cardinality which is loosely based on the inception module as shown in Fig. 12. The cardinality refers to the number of independent and identical paths, where each path performs transformation of the input data, divided along the channel dimension Xie et al. (2017). In other words, instead of solely relying on increasing the depth of the model, ResNext enhances the feature learning by parallelizing the feature extraction through this cardinal path. In the proposed architecture, each path configuration is similar to the residual block of ResNet. The output from each path is then aggregated to form a comprehensive and diverse representation of the input data. The skip connection is used to mitigate the vanishing gradient problem. WideResNet is a CNN architecture that is based on ResNet with the aim to mitigate diminishing feature reuse problem. Instead of making the network deeper, WideResNet makes the network wider by increasing the number of channels by $k$ factor Zagoruyko and Komodakis (2017). The increased width allows the model to capture a more diverse features, enhancing its ability to learn complex relationships in the input data.
Figure 12: A cardinal block.
3.2 Deep Unsupervised Learning
Deep unsupervised models are trained with an unlabelled dataset. The learning process of these models involve calculating the prediction error through a loss function and utilizing the error to adjust the weights iteratively until the prediction error is minimized. Among the deep supervised models, three important models are identified namely multilayer perceptron, convolutional neural network and recurrent neural network.
Restricted Boltzmann Machine is a generative neural network model that learns a probability distribution based on a set of inputs. The model consists of a visible (input) layer and a hidden layer with symmetrically weighted connections as shown in Fig. 13. The input layer represents the input data with each node corresponding to a feature or variable while the hidden layer learns the abstract representation of the input data. Restricted Boltzmann machine model is trained using contrastive divergence, an algorithm that is based on a modified form of gradient descent, utilizing a sampling-based approach to estimate the gradient Hinton (2012). It has found success in solving combinative problems such as dimensionality reduction, collaborative filtering and topic modelling.
Figure 13: A restricted Boltzmann machine.
Deep Belief Network can be viewed as a stack of restricted Boltzmann machines, comprising a visible layer and multiple hidden layers Hinton et al. (2006) as shown in Fig. 14. Deep belief network has two training phases. The initial phase is known as pretraining in which the network is trained layer by layer, with each layer serves as a pretraining layer of the subsequent layers. This sequential learning allows the hidden layers learn complex hierarchical feature representation of the data. The second phase is called fine-tuning whereby the deep belief network model can be further trained with supervision to perform tasks such as classification and regression Hinton (2009).
Figure 14: A deep belief network.
Autoencoder is a generative neural network model that learns to encode the input data into a compressed representation and then reconstructs the original data from this representation. The layers that encode the input data is known as encoder while the layers that responsible for the reconstruction is referred to as the decoder as shown in Fig. 15. The encoded data (hidden layer) represents the abstract features of the input data also known as latent space or encoding. The decoder can be removed from the autoencoder, creating a standalone model that can be used for data compression and dimensionality reduction Romero et al. (2017), Li et al. (2020c). The decoder can also be replaced with predictive layers for classification task Mohd Noor (2021). The network architecture can be built using fully-connected and convolutional layers Li et al. (2023).
Figure 15: An autoencoder.
Several autoencoder variants have been introduced to improve the autoencoder’s ability to capture better feature representation. Some introduced penalty terms to the loss function such as sparsity penalty (sparse autoencoder) Ng and others (2011) to encourage sparse representation and Jacobian Frobenius norm (contractive autoencoder) Rifai et al. (2011) to be less sensitive to small and insignificant variations in the input data while encoding the feature representation. Others trained the autoencoder to recover original data from corrupted data with noise Vincent et al. (2008). An improved denoising autoencoder knowns marginalized denoising autoencoder has been proposed which marginalizes the noise by adding a term that is linked to the encoding layer Chen et al. (2012). Variational autoencoder is a variant of autoencoder that has similar architecture i.e. encoder, latent space and decoder. Despite the similarity, instead of learning a fixed encoding, variational autoencoder learns the probability distribution of the input data in the latent space Kingma and Welling (2013). The model can be used to generate data by sampling from the learned probability distribution. The network architecture can be built by stacking more than one fully-connected layer and convolutional layer.
Generative Adversarial Network (GAN) is another generative neural network model that is designed for generating data that adheres closely to the distribution of the original training set. The model consists of two different neural networks namely generator and discriminator as shown in Fig. 16. The generator learns to imitate the distribution of the training set given a noise vector, effectively outsmarting the discriminator. Simultaneously during the training, the discriminator is trained to differentiate between the real data from the training set and synthetic data generated by the generator Goodfellow et al. (2014). This intricate dynamic between the networks drives an iterative learning process whereby the generator continually refines its ability to create synthetic data that closely resembles the real data while the discriminator enhances its ability to distinguish between authentic and fake data.
Figure 16: A generative adversarial network.
The model can be extended by providing the labels to both generator and discriminator in which the model known as conditional GAN, capable of generating 1000 image classes Odena et al. (2017). Conditional GANs require a labelled dataset which might limit its application. InfoGAN is similar to conditional GAN, but the labels are substituted with latent codes, which allows the model to be trained in an unsupervised manner Chen et al. (2016). GANs often suffer from mode collapse whereby the model can only generate a single or small set of outputs. Wasserstein GAN improves the training by utilizing Wasserstein loss function which measures the difference between the real and synthesized data distribution Weng (2019). ProGAN tackles the training instability of GAN by progressively growing the generator and discriminator. The idea is that the model is scaled up gradually, starting with the simplest form of the problem and little by little the problem’s complexity is increased as the training progresses Karras et al. (2018). StyleGAN leverages the progressive GAN’s approach and neural style transfer to improve quality of the generated data Karras et al. (2019). The model is characterized by the independent manipulation of both style and content, allowing it to generate diverse styles and high quality data.
4 Applications of Deep Learning
As discussed in the previous section, the application of deep learning ranges from computer vision Tan et al. (2020a), natural language processing Otter et al. (2021), healthcare Esteva et al. (2019), robotics Soori et al. (2023), education Hernández-Blanco et al. (2019), and many others. This section presents the applications of deep learning across several areas.
4.1 Computer Vision
Computer vision is an essential field in artificial intelligence. It is a field of study that focuses on enabling computers to acquire, analyze and interpret visual inputs to derive meaningful information. The visual inputs can take many forms such as digital images, sequence of images or video and point cloud, and the source of these inputs can be camera, LiDaR, medical scanning machine etc. Deep learning, specifically CNN models have been widely used in real-world computer vision applications including image classification, object detection and image segmentation. This section discusses more details about the recent advancements in deep learning models that have been achieved over the past few years.
4.1.1 Image Classification
Image classification is a fundamental task in computer vision which involves categorizing an image into one of predefined classes based on the visual content. The objective of image classification is to enable computers or machines to differentiate between objects within images, in a manner similar to how humans interpret visual information. Image classification is a crucial component in various applications such as robotics, manufacturing and healthcare. LeNet-5, introduced in 1998, is one of the earliest convolutional neural networks that was successfully trained to classify handwritten digits. The model underwent a series of improvements, including the use of tanh and average pooling which enhanced its ability to extract hierarchical features, ultimately improving overall performance. The model architecture comprises two convolutional layers, each with an average pooling layer, followed by two fully-connected layers, including the output layer Lecun et al. (1998). Since then, numerous CNN models have been proposed based on LeNet-5 for image classification Simard et al. (2003); Matsugu et al. (2003) but the most significant one is AlexNet in 2012 which saw a transformative breakthrough in deep learning. AlexNet is considered the first CNN model with a large number of parameters that significantly improved the performance of image classification on a very large dataset (ImageNet). The model won first place in ILSVRC 2012, improving the test error from the previous year by almost 10% Krizhevsky et al. (2012). Numerous significant CNN models have been introduced in subsequent ILSVRC competitions including ZFNet, VGG16, GoogleNet, ResNet and ResNext. In general, the research focused on increasing the number of layers, addressing the problem of vanishing gradient and diminishing of feature reuse.
Research in image classification continues to evolve with a focus on addressing key challenges to improve the classification performance. One notable trend is the formulation of the loss function to address problems such as neglecting well-classified instances and imbalance distribution of class labels. In a particular study, an additive term is introduced to the cross-entropy loss to reward the models for the correctly classified instances. This formulation encourages the models to also pay attention to well-classified instances while focusing on the bad-classified ones Zhao et al. (2022). Another study proposes an asymmetric polynomial loss function using the Taylor series expansion. The loss function allows the training to selectively prioritize contributions of positive instances to mitigate the issue of imbalance between negative and positive classes Huang et al. (2023b). The asymmetric polynomial loss requires a large number of parameters to be fine-tuned and may lead to overfitting. A robust asymmetric loss is formulated by introducing a multiplicative term to control the contribution of the negative gradient and making it less sensitive to parameter optimization Park et al. (2023). Combining multiple deep learning models improves the overall performance by leveraging the diverse strengths of individual models. However, identifying the optimal combination is non-trivial due to the large number of hyperparameters. A straightforward method is to employ the weighted sum rule Nanni et al. (2023). To enhance the overall performance, an algorithm, named greedy soups, adds a model based on the validation accuracy Wortsman et al. (2022). The final prediction is produced via averaging. Multi-symmetry ensembles framework improves the building of diverse deep learning models by utilizing contrastive learning Loh et al. (2023). Then, the diverse models are sequentially combined based on their validation accuracy.
Vision transformers (ViT) offers an alternative to convolutional neural networks that have long been the dominant architecture for image classification, by leveraging self-attention mechanisms for scalable representation learning. Despite its effectiveness, ViT is sensitive to hyperparameter optimization and substandard performance on smaller datasets Xiao et al. (2021). Furthermore, ViT lacks the ability to leverage local spatial features which is inherent in convolutional neural networks Wu et al. (2021). Therefore, several studies attempt to incorporate convolutional layers into ViT architecture to improve its performance and robustness. In particular, conformer is a network architecture with two branches: CNN branch and a transformer branch to extract local and global features respectively Peng et al. (2023). Both branches are connected by two “bridges” of $1\times 1$ convolution and up or down sampling operations, allowing the branches to share their features and enhance the feature representation. Both branches output predictions which are combined to produce the final prediction. A hybrid architecture, named MaxViT, combines convolutional networks and vision transformer to address the lack of scalability issues of self-attention mechanisms when the model is trained on large input size Tu et al. (2022). The improved vision transformer is composed of two modules whereby the first module attends local features in non-overlapping image patches and the global features are attended by processing a grid of sparse and uniform pixels. The transformer is stacked with a block of convolutional layers to extract local spatial features. The architecture of MaxViT is shown in Figure 17. Another study proposes a convolutional transformer network, introducing the depthwise convolutional block into the ViT Ma et al. (2024). This configuration allows the model to exploit the ability of convolutional networks to extract local spatial features while the ViT attends the extracted local features to focus on relevant information, enhancing the model’s ability to capture complex patterns and relationships.
Figure 17: The architecture of MaxViT Tu et al. (2022).
4.1.2 Object Detection
Deep learning plays a major role in significantly advancing the state-of-the-art in object detection performance. Region-based CNN (R-CNN) is the first breakthrough in object detection that combines CNN with selective region proposals Girshick et al. (2014). The region proposals are the candidate bounding boxes serving as the potential region of interests (objects) within the input image, and the CNN are used to extract features from the region proposals and classify the regions for object detection. An improved model, named Fast R-CNN introduces two prediction branches: object classification and bounding box regression which improves the overall performance of object detection Girshick (2015). However, R-CNN and Fast R-CNN models are computationally expensive and slow, thus practically infeasible for real-time applications. Addressing this issue, Fast R-CNN is integrated with a region proposal network, referred to as Faster R-CNN Ren et al. (2015). The region proposal network (RPN) is used to efficiently generate region proposals for object detection. RPN takes an input image and output a set of rectangle object proposals, each with a confidence score to indicate the likelihood of an object’s presence. To this end, RPN introduces the concept of anchor boxes whereby multiple bounding boxes of different aspect ratios are defined over the feature maps produced by the convolutional networks. These anchor boxes are then regressed over the feature maps to localize the objects, contributing to the improved speed and effectiveness of Faster R-CNN. The training of Faster R-CNN is divided into two stages. First the RPN is pre-trained to generate the region proposals and then, the Fast R-CNN is trained using the region proposals generated by the RPN for object detection. The backbone network responsible for extracting the features for Faster R-CNN is either ZFNet or VGG16.
In two-stage object detectors, the region proposals are generated first, and then used for object detection. The two-stage process is computationally intensive and infeasible for real-time object detection applications. You Only Look Once or YOLO proposes a one-stage detection by directly predicting bounding boxes and object’s confidence score in a single forward pass through the neural network Redmon et al. (2016). This single pass architecture significantly reduces the computational complexity, making YOLO suitable for real-time object detection applications. In YOLO, the input image is divided into S×S grids, each grid cell is responsible for detecting the objects present in the cell. Specifically, each grid cell predicts multiple bounding boxes and associated object’s confidence score, enabling simultaneous object detection across the entire image. Subsequent enhancements such as YOLOv3 Redmon and Farhadi (2018) and YOLOv4 Bochkovskiy et al. (2020) are proposed, improving the model’s capability and accuracy. Single Shot Multibox Detector (SSD) is another one-stage detector, aims to address the issue of real-time object detection Liu et al. (2016). SSD also eliminates the region proposal generation and directly predicts bounding boxes and confidence scores, reducing the computational complexity. To improve the overall performance, SSD produces the predictions from different levels of feature maps, allowing detection of objects of different sizes in the input image.
A common issue in object detection problems is the extremely imbalanced ratio of foreground to background classes. Addressing this issue, RetinaNet introduces a loss function that is based on the cross entropy called focal loss. Focal loss reduces the loss contribution of easily classified objects, allowing the model training to focus on the difficult objects Lin et al. (2020). RetinaNet adopts the Feature Pyramid Network (FPN) Lin et al. (2017a) with ResNet as the backbone network for extracting the feature maps. FPN is a network architecture with a pyramid structure that efficiently captures multiscale feature representation, facilitating object detection across various sizes. To further improve the overall performance, EfficientDet introduces bi-directional FPN which incorporates multi-level feature fusion to better capture multiscale feature representation Tan et al. (2020b). Also, the model utilizes EfficientNet Tan and Le (2019) as the backbone network to achieve a balance between computational efficiency and accuracy.
Object detection performance relies on a post-processing step called non-maximum suppression (NMS) to eliminate duplicate detections and select the most relevant bounding boxes. Specifically, NMS sorts all detection boxes based on their confidence scores, selects a box with the maximum score and discards the other boxes with a significant overlap with the selected box. This process is repeated on the remaining detection boxes. However, due to the inconsistency between the confidence score and the quality of object localization, NMS retains poorly localized bounding boxes with high confidence score while discarding more accurate predictions with poor confidence score. To mitigate this limitation, instead of discarding the neighboring boxes with significant overlap, soft-NMS applies Gaussian function to lower their confidence scores Bodla et al. (2017). The idea is not to discard the neighboring bounding boxes, but gradually decline their scores based on the extend of the overlap with the selected box. This results in a smoother suppression, preserving the better-localized bounding boxes. Adaptive NMS introduces an adaptive threshold for the suppression of bounding boxes Liu et al. (2019a). The algorithm dynamically adjusts the threshold based on the level of overlapping of the selected box with the other bounding boxes.
Detection Transformer (DeTR) is an end-to-end trainable object detection model that leverages the transformer architecture to eliminates the need for handcrafted components such as anchor boxes and non-maximum suppression Carion et al. (2020). The self-attention mechanism of the transformer captures the global context and relationships between different parts of the image, allowing it to localize the objects and remove duplicate predictions. The model is trained with a set of loss functions that perform bipartite matching between the predicted and ground truth objects. DeTR uses ResNet as backbone network. Despite the success of DeTR in simplifying and improving object detection tasks, DeTR suffers from a long training time and low performance at detecting small objects due to its reliance on self-attention mechanism of the transformer, which lacks a multiscale feature representation. To mitigate this limitation, Deformable DeTR introduces a multiscale deformable attention module which can effectively capture feature representation at different scales Zhu et al. (2020). Furthermore, the attention module leverages deformable convolution, allowing the model to adapt to spatial variation and capture more informative features in the input data. Dynamic DeTR addresses the same issues by utilizing a deformable convolution-based FPN to learn multiscale feature representation Dai et al. (2021). Moreover, the model replaces the transformer encoder with a convolution-based encoder to attend to various spatial features and channels. This modification allows the model to effectively detect small objects and converge faster during training. The architecture of dynamic DeTR is shown in Figure 18. A training scheme known as Teach-DeTR is proposed to improve the overall performance of DeTR Huang et al. (2023a). The training scheme leverages the predicted bounding boxes by other object detection models during the training by calculating the loss of one-to-one matching between the object queries and the predicted boxes.
Figure 18: The architecture of dynamic DeTR Dai et al. (2021).
4.1.3 Image Segmentation
Image segmentation is another important task in which deep learning has a significant impact. One of the earliest deep learning models for image segmentation is the fully convolutional network Long et al. (2015). A fully convolutional network consists of only convolutional layers which accepts an input of an arbitrary size and produce the predicted segmentation map of the same size. The authors adopt the AlexNet, VGG16 and GoogleNet, replace their fully connected layers with convolutional layers and append a 1×1 convolutional layer, followed by bilinear up-sampling to match the size of the input. The model was considered a significant milestone in image segmentation, demonstrating the feasibility of deep learning for semantic segmentation trained in end-to-end manner. Deconvolution network is another popular deep learning model for semantic segmentation Noh et al. (2015). The model architecture consists of two parts: encoder and decoder. The encoder takes an input image and uses the convolutional layers to generate the feature maps. The feature maps are fed to the decoder composed of un-sampling and deconvolutional layers to predict the segmentation map. SegNet is another encoder-decoder model for semantic segmentation Badrinarayanan et al. (2017). The encoder is a sequence of convolutional (with ReLU) and max-pooling blocks which is analogous to a convolutional neural network. The decoder is composed of up-sampling layers which up-samples the inputs using the memorized pooled indices generated in the encoder phase, and convolutional layers without non-linearity. The encoder progressively reduces the resolution of the input data while extracting abstract features through a series of convolutional and pooling layers. This process causes the loss of fine-grained information, degrading the overall performance of segmentation. LinkNet mitigates this limitation by passing the feature maps at several stages generated by the encoder to the decoder, hence reducing information loss Chaurasia and Culurciello (2017). The model architecture of LinkNet is similar to SegNet, but utilizes ResNet as the encoder.
While Faster R-CNN is a significant approach in object detection task, it has been extended to perform instance segmentation task. One such extension is Mask R-CNN which is based on Faster R-CNN, introduces an additional branch for predicting the segmentation mask He et al. (2017). Similar to Faster R-CNN, Mask R-CNN utilizes the RPN to generate region proposals and then the region of interest alignment is applied to extract more accurate features from the proposed regions. Mask R-CNN does not leverage the multiscale feature representation which may degrade the overall performance of segmentation. To overcome this limitation, Path Aggregation Network (PANet) incorporates the FPN and introduces a bottom-up pathway to facilitate the propagation of the low-level information Liu et al. (2018). The pathway takes the feature maps of the previous stage as input and performs 3×3 convolution with stride 2 to reduce the spatial size of the feature maps. The generated feature maps are then fused with the feature maps from the FPN through the lateral connection. The model adopts the three branches as in Mask R-CNN. MaskLab is an instance segmentation model based on Faster R-CNN, consisting of object detection, segmentation and instance (object) center direction prediction branches Chen et al. (2018). The direction prediction provides useful information to distinguish instances of the same semantic label, allowing the model to further refine the instance segmentation results.
Attention mechanisms have been integrated into the segmentation models to learn the weights of multiscale features at each pixel location. A multistage context refinement network introduces a context attention refinement module that is composed of two parts, context feature extraction and context feature refinement Liu et al. (2023a). The context feature extraction captures both local and global context information, fuses both contextual information and passes it to the context feature refinement while the context feature refinement removes redundant information and generates a refined feature representation, improving the utilization of contextual information. The context attention is added to the skip connection between the encoder and the decoder. Handcrafted features are often abandoned for automatic feature extraction using convolutional networks. However, it is argued that the interpretability and domain-specific knowledge embedded in handcrafted features can provide valuable insights. To this end, an attention module based on the covariance statistic is introduced to model the dependencies between local and global context of the input image Liu et al. (2022). Two types of attention are introduced: spatial covariance attention focuses on the spatial distribution and channel covariance attention attends to the important channels. Furthermore, the covariance attention does not require feature shape conversion, hence significantly reducing the space and time complexity of the model.
The convolutional layers use local receptive fields to process input data which can be effective for exploiting spatial patterns and hierarchical features but may find it difficult to capture global relationships across the entire image. The ViT has been leveraged to mitigate this issue in semantic segmentation Strudel et al. (2021). Specifically, the input image is divided into patches and treated as input to the transformer to capture the global relationship between the patches, significantly improving the prediction of the segmentation map. Global context ViT aims to address the lack of ViT’s ability to leverage local spatial features Hatamizadeh et al. (2023). As shown in Figure 19, the transformer consists of local and global self-attention modules. The role of global self-attention is to capture the global contextual information from different image regions while the short-range information is captured by the local self-attention. Multiscale feature representation is crucial for accurate semantic segmentation. However, the transformer often combines the features without considering their appropriate (optimal) scales, thus affecting the segmentation accuracy. Transformer scale gate is a module proposed to address the issue of selecting an appropriate scale based on the correlation between patch-query pairs Huang et al. (2023a). The transformer takes attention (correlation) maps as input and calculates the weights of the multi-scale features for each image patch, allowing the model to adaptively choose the optimal scale for each patch.
Figure 19: The architecture of global context ViT Hatamizadeh et al. (2023).
4.1.4 Image Generation
Image generation refers to the process of creating images based on input texts. Generally, the task can be divided into three stages. The first stage is extracting features from the input text, followed by generating the image and finally controlling the image generation process to ensure the output meets specific criteria and constraints. This section focuses on the progress made in the development of deep learning models of the second stage (image generation) since it directly impacts the quality of the generated images. Variational autoencoder is one of the earliest deep learning models that is capable of generating images Kingma and Welling (2013). Variational autoencoder learns to generate data by capturing the underlying (Gaussian) distribution of the training data. During the generation process, the distribution parameters are sampled and passed to the decoder to generate the output image. Although the generated images are blurry and unsatisfactory, it has shown a lot of potential in image generation tasks. The introduction of GAN significantly improved the quality of generated images. GAN consists of two connected neural networks, a generator and a discriminator that are trained simultaneously in a competitive manner Goodfellow et al. (2014). The generator learns to generate realistic images to fool the discriminator, while the discriminator learns to distinguish between fake and real images. The generated images are less blurry and more realistic. Several enhanced models have been proposed to improve its usability and overall performance such as CGAN Odena et al. (2017) which allows us to tell what image to be generated, and the deep convolutional GAN (DCGAN) Radford et al. (2015) which provides a more stable structure for image generation. DCGAN is the basis of many subsequent improvements in GANs.
StackGAN divides the process of image generation into two stages Zhang et al. (2017). Stage-I generates a low-resolution image by creating basic shapes and colors and the background layout using the random noise vector. Stage-II completes the details of the image and produces a high-resolution photo-realistic image. StackGAN++ is the enhanced model of StackGAN whereby it consists of multiple generators with shared parameters to generate multiscale images Zhang et al. (2018a). The generators have a progressive goal with the intermediate generators generating images of varying sizes and the deepest generator producing the photo-realistic image. HDGAN is a generative model featuring a single-stream generator with hierarchically nested discriminators at intermediate layers Zhang et al. (2018b). These layers, each connected to a discriminator, generate multiscale images. The lower resolution outputs are used to learn semantic image structures while the higher resolution outputs are used to learn fine-grained details of the image. StackGAN heavily relies on the quality of the generated image in Stage-I. DM-GAN incorporates a memory network for image refinement to cope with badly generated images in Stage-I Zhu et al. (2019b). The memory network dynamically selects the words that are relevant to the generated image, and then refines the details to produce better photo-realistic images.
AttnGAN is the first to incorporate attention mechanisms into the multiple generators to focus on words that are relevant to the generated image Xu et al. (2018). To this end, in addition to encoding the whole sentence into a global sentence vector, the text encoder encodes each word into a word vector as shown in Figure 20. Then, the image vector is used to attend to the word vector using the attention modules at each stage of the multistage generators. Furthermore, AttnGAN introduces a loss function to compute the similarity between the generated image and the associated sentence, improving the performance of image generation. A similar work is reported whereby the model known as ResFPA-GAN, incorporates attention modules into the multiple generators Sun et al. (2019). Specifically, a feature pyramid attention module is proposed to capture high semantic information and fuse the multiscale feature, enhancing the overall performance of the model. DualAttn-GAN improves AttnGAN by incorporating visual attention modules to focus on important features along both spatial and channel dimensions Cai et al. (2019). This allows the model to better understand and capture both the context of the input sentence and the fine details of the image, resulting in more realistic image generation.
Figure 20: The architecture of AttnGAN Xu et al. (2018).
Although multistage generators improve image generation performance by leveraging multiscale representation, the generated images may contain fuzzy shapes with coarse features. DF-GAN replaces the multistage generators with a single-stage deep generator featuring residual connections and trained with hinge loss Tao et al. (2022). Furthermore, DF-GAN introduces a regularization strategy on the discriminator that applies a gradient penalty on real images with matching text, allowing the model to generate more text-matching images. DMF-GAN an improved DF-GAN, incorporates three novel components designed to leverage semantic coherence between the input text and the generated image Yang et al. (2024). The first component is the recurrent semantic fusion module, which models long range dependencies between the fusion blocks. The second component is the multi-head attention module which is placed towards the end of the generator to leverage the word features, forcing the generator to generate images conditioned on the relevant words. The last component is the word-level discriminator which provides fine-grained feedback to the generator, facilitating the learning process and improving the overall quality of the generated images. Figure 21 shows the architecture of DMF-GAN. The process of image generation involves feeding a noise vector to the generator at the very beginning of the network. However, as the generator goes deeper, the noise effect may be diminished, affecting the diversity of the image generation results. To mitigate this issue, DE-GAN incorporates a dual injection module into the single-stage generator Jiang et al. (2024). The dual injection module consists of two text fusion layers followed by a noise broadcast operation. The text fusion layer takes the sentence embedding and fuses it with the input feature map using the fully-connected layer. Then noise is injected into the output feature map to retain the randomness in the generation process, improving diversity and generalization of the model.
Figure 21: The architecture of DMF-GAN Yang et al. (2024).
4.2 Time Series and Pervasive Computing
Pervasive computing, often referred to as ubiquitous computing, is the process of integrating computer technology into everyday objects and surroundings so that they become intelligent, networked, and able to communicate with one another to offer improved services and functionalities Weiser (1991). According to He et al He et al. (2020b), the role of pervasive computing is foremost in the field where it provides the ability to distribute computational services to the surroundings where people work, leading to trust, privacy, and identity. Examples of pervasive computing applications include smart homes with connected appliances, wearable devices that monitor health and fitness, smart cities with sensor networks for traffic management, and industrial applications that utilize the Internet of Things (IoT) for monitoring and control. Generally, the continuous interaction of interconnected devices in pervasive computing often result in time series data, which captures the evolution of various parameters over time.
For instance, medical sensors, such as electrocardiograms (ECG) and electroencephalograms (EEG), generate time series data that contain critical diagnostic information, which deep learning can use to detect anomalies, predict diseases, and classify medical conditions with improved accuracy. Also, devices such as accelerometers, magnetometers and gyroscopes, among others can be used to capture human activity signals, which are often represented as time series of state changes Ige and Noor (2022). In traditional machine learning, features such as mean, variance and others are manually extracted from times series of state changes before human activity classification. However, deep learning models automatically extract features Mohd Noor (2021). Also, in other fields such as finance, which entail time series data, deep learning has been instrumental in stock price prediction Singh and Srivastava (2017), fraud detection Zhang et al. (2021), and algorithmic trading Lei et al. (2020), among others. Generally, deep learning networks excel at capturing intricate temporal relationships within time-series data, enabling more precise predictions and improved decision-making. Based on this, several deep learning models have been employed for feature learning across various time series and pervasive computing domains.
4.2.1 Human Activity Recognition
Human activity recognition (HAR) finds application across various domains including intelligent video surveillance, environmental home monitoring, video storage and retrieval, intelligent human-machine interfaces, and identity recognition, among many others. It includes various research fields, including the detection of humans in video, estimating human poses, tracking humans, and analyzing and understanding time series data Zhang et al. (2019a). Despite the advancements in vision-based HAR, there exist inherent limitations. Generally, vision-based approaches heavily rely on camera systems, which may have restricted views or be affected by lighting conditions, occlusions, and complex backgrounds Ige and Noor (2022). Additionally, vision-based HAR struggles with identifying actions that occur beyond the range of the camera or actions that are visually similar.
Wearable sensors offer a promising alternative to overcome these limitations. By directly capturing data from the individual, wearable sensors provide more comprehensive and accurate information about human activities. The signals obtained from wearable sensors typically represent time series data reflecting state changes in activities. Deep learning models can effectively learn from these signals, allowing for robust and accurate recognition of human activities. Moreover, wearable sensors offer the advantage of mobility, enabling activity recognition in various environments and situations where vision-based systems may be impractical or ineffective Dang et al. (2020). Generally, the time series nature of signals from wearable sensors presents an excellent opportunity for deep learning models to excel in recognizing human activities with high accuracy and reliability.
Several researchers have proposed the use of CNN, RNN, and Hybrid models for deep learning based feature learning in wearable sensor HAR. For instance, using two-dimensional CNN (Conv2D), several researchers, as seen in Gao et al. (2021a), Gupta (2021) and Erdaş and Güney (2021), among others, have developed deep learning models for wearable sensor HAR, despite the time series nature of the data. This is often done by treating the time series signals from wearable sensors as 2D images by reshaping them appropriately. To achieve this, researchers often organize each time series signal into a matrix format, with time along one axis and sensor dimensions along the other, before creating a pseudo-image representation, which allows the matrix to be be fed into Conv2D layers for feature extraction. Conv2D layers excel at capturing spatial patterns and relationships within images, and by treating the time series data as images, these layers can learn relevant spatial features that contribute to activity recognition. The convolution operation performed by Conv2D filters across both the time and sensor dimensions, allowing the network to identify patterns and features that may be indicative of specific activities.
Even though Conv2D can effectively capture spatial dependencies within the data, it often struggles to capture temporal dependencies inherent in time series data. Since Conv2D processes data in a grid-like fashion, it does not fully leverage the sequential nature of the time series, potentially leading to less effective feature extraction for wearable sensor HAR tasks. For this reason, recent HAR architectures have leveraged one-dimensional CNN (Conv1D) and other RNNs for automatic feature extraction. Conv1D layers are specifically designed to capture temporal dependencies within sequential data. They operate directly on the time series data without reshaping it into a 2D format, allowing them to capture temporal patterns more effectively. Conv1D layers are better suited for extracting features from time series data, making them a more natural choice for wearable sensor HAR Mohd Noor (2021).
For instance, Ragab et al. Ragab et al. (2020) proposed a random search Conv1D model, and evaluated the performance of the model on UCI-HAR dataset. The result showed that the model achieved a recognition accuracy of 95.40% when classifying the six activities in the dataset. However, the model exhibited extended training times due to the dynamic nature of some activities within the dataset. To address this, Banjarey et al. (2022) proposed the use of varying kernel sizes in Conv1D layers to recognize various activities, including sitting, standing, walking, sleeping, reading, and tilting. Also, a few Conv1D layers were stacked in order to streamline the time optimization process for training the neural network. Also, some researchers have proposed models that combine machine learning algorithms with Conv1D in HAR, as seen in Shuvo et al. Shuvo et al. (2020). Their work presented a two-stage learning process to improve HAR by classifying activities into static and dynamic using Random Forest, before using Support Vector Machine to identify each static activity, and Conv1D to recognize dynamic activities. The result showed that the method achieved an accuracy of 97.71% on the UCI-HAR dataset.
Following these advancements, several researchers have further explored Conv1D architectures with various modifications, to enhance feature learning in activity recognition systems. For example, Han et al. Han et al. (2022) developed a two-stream CNN architecture as a plug-and-play module to encode contextual information of sensor time series from different receptive field sizes. The module was integrated into existing deep models for HAR at no extra computation cost. Experiments on OPPORTUNITY, PAMAP2, UCI-HAR and USC-HAD datasets showed that the module improved feature learning capabilities. A similar research reported in Ige and Noor (2023) proposed the WSense module to address the issue of differences in the quality of features learnt, regardless of the size of the sliding window segmentation, and experimented on PAMAP2 and WISDM datasets. The results showed that by plugging the WSense module into Conv1D architectures, improved activity features can be learned from wearable sensor data for human activity recognition.
Similarly, some researchers have also proposed the use of standalone RNNs in HAR, and a hybrid of Conv1D architectures with RNNs such as LSTMs Deep and Zheng (2019), BiLSTMs Luwe et al. (2022); Shi et al. (2023), GRUs Dua et al. (2023) and BiGRUs Imran et al. (2023); Chen et al. (2022b) in order to fully harness the feature learning capabilities of both CNN and RNNs. For instance, Nafea et al. Nafea et al. (2021), leveraged Bi-LSTM and Conv1D with increasing kernel sizes to learn features at various resolutions. Human activity features were extracted using the stacked convolutional layers with a Bi-LSTM layer, before including a flattening layer and a fully connected layer for subsequent classification. However, the model had issues extracting quality features of dynamic activities compared to static activities. To address such issues, some research works have incorporated attention mechanisms in Conv1D-based architectures to improve feature learning of dynamic and complex activities from time series signals obtained from wearable sensors. For example, Khan and Ahmad Khan and Ahmad (2021) designed three lightweight convolutional heads, with each specialized in feature extraction from wearable sensor data. Each head comprised stacked layers of Conv1Ds, along with embedded attention mechanisms to augment feature learning. The results demonstrated that integrating multiple 1D-CNN heads with attention mechanisms can enhance feature learning for Human Activity Recognition (HAR). These diverse modifications and adaptations showcase the versatility and potential of deep learning models in achieving state-of-the-art in HAR systems.
Figure 22: The architecture of multi-head CNN model Khan and Ahmad (2021).
4.2.2 Speech Recognition
Speech, as the primary mode of human communication, has captivated researchers for over five decades, especially since the inception of artificial intelligence Nassif et al. (2019). From the earliest endeavors to understand and replicate the complexities of human speech, to contemporary advancements leveraging cutting-edge technologies, the quest for accurate and efficient speech recognition systems has been relentless. In recent years, the emergence of deep learning techniques has revolutionized the speech recognition field. Deep learning has demonstrated unparalleled success in processing and extracting intricate patterns from vast amount of data. When applied to the realm of speech recognition, deep learning have surpassed traditional approaches by learning intricate features directly from raw audio signals, circumventing the need for handcrafted features and complex preprocessing pipelines. This paradigm shift has significantly advanced the state-of-the-art in speech recognition, enabling systems to achieve unprecedented levels of accuracy and robustness across various languages, accents, and environmental conditions. Generally, deep learning has been extended to other essential applications of speech recognition, such as speaker identification Tirumala and Shahamiri (2016); Ye and Yang (2021), emotion recognition Khalil et al. (2019), language identification Singh et al. (2021), accent recognition Jiao et al. (2016), age recognition Sánchez-Hevia et al. (2022) and gender recognition Alnuaim et al. (2022), among many others.
Prior to the adoption of deep learning in speech recognition, the foundation of traditional speech recognition systems was the use of Gaussian Mixture Models (GMMs), which are often combined with Hidden Markov Models (HMMs) to represent speech signals Srivastava and Pandey (2022). This is because a speech signal can be thought of as a short-term stationary signal. The spectral representation of the sound wave is modelled by each HMM using a mixture of Gaussian. However, they are considered statistically inefficient for modelling non-linear or near non-linear functions Padmanabhan and Premkumar (2015); Nassif et al. (2019). This is because HMMs rely on a set of predefined states and transition probabilities, making assumptions about the linearity and stationarity of the underlying data. While suitable for modelling certain aspects of speech, HMMs often fall short when tasked with representing the intricate nonlinearities and variability present in speech signals. Speech, by nature, exhibits nonlinear and dynamic characteristics, with features such as intonation, rhythm, and phonetic variations challenging the simplistic assumptions of traditional statistical models like HMMs. In other words, GMM-HMM approach had limitations in capturing complex acoustic patterns and long-term dependencies in speech Mukhamadiyev et al. (2022).
In recent times, CNN and RNNs have been leveraged for automatic speech recognition in order to consider a longer or variable temporal window for context information extraction Lu et al. (2020). Generally, CNNs are well-suited for capturing local patterns and hierarchical features in data, making them effective for modelling acoustic features in speech. By directly learning features from raw speech signals, CNNs bypassed the need for handcrafted features used in traditional GMM-HMM systems. Additionally, CNNs can capture long-range dependencies in the data, which is crucial for understanding the context of speech. Likewise, the RNNs are suitable choice for exploring extended temporal context information in one processing level for feature extraction and modelling.
Based on this, several researchers have proposed the use of both CNN and variants of RNNs for automatic speech recognition and for other speech related tasks. For instance, Hema and Garcia Marquez (2023), used CNN to classify speech emotions and benchmarked on a dataset consisting of seven classes (anger, disgust, fear, happiness, neutral, sadness and surprise). However, CNN lack the ability to model temporal dependencies explicitly. In speech recognition, understanding the temporal context of speech is essential for accurate transcription. Also, speech signals are inherently sequential, and information from previous time steps is crucial for understanding the current speech segment. CNNs, by design, do not inherently capture this sequential nature. For this reason, variants of RNNs have been leveraged to collect extended contexts in speeches. This is because RNNs are designed to model sequential data by maintaining hidden states that capture information from previous time steps. This allows them to capture temporal dependencies effectively, making them well-suited for ASR tasks. In Shewalkar et al. (2019), the authors evaluated the performance of RNN, LSTM, and GRU on a popular benchmark speech dataset (ED-LIUM). The results showed that LSTM achieved the best word error rate while the GRU optimization was faster and achieved word error rate close to that of LSTM.
However, RNN architectures process input sequences sequentially, which limits their ability to capture global context information effectively. As a result, they may struggle to understand the entire context of a spoken utterance, leading to lower transcription accuracy, particularly in tasks requiring understanding beyond local dependencies. Also, most CNN and RNN automatic speech recognition systems comprise of separate acoustic, pronunciation, and language modelling components that are trained independently. Usually, the acoustic model bootstraps from an existing model that is used for alignment in order to train it to recognise context dependent (CD) states or phonemes. The pronunciation model, curated by expert linguists, maps the sequences of phonemes produced by the acoustic model into word sequences. For this reason, Sequence-to-Sequence (Seq2Seq) models are being proposed in automatic speech recognition to train the acoustic, pronunciation, and language modelling components jointly in a single system Prabhavalkar et al. (2017). Seq2Seq methods in automatic speech recognition are a class of models that aim to directly transcribe an input sequence of acoustic features such as speech spectrograms or Mel-frequency cepstral coefficients into a sequence of characters or words representing the recognized speech. There have been a variety of sequence-to-sequence models explored in the literature, including Recurrent Neural Network Transducer (RNN-T) Graves (2012), Listen, Attend and Spell (LAS) Chan et al. (2015), Neural Transducer Jaitly et al. (2016), Monotonic Alignments Raffel et al. (2017) and Recurrent Neural Aligner (RNA) Sak et al. (2017).
Figure 23: Sequence-to-Sequence
As shown in Figure 23, the encoder component takes the input sequence of acoustic features and processes it to create a fixed-dimensional representation, often called the context vector. This representation captures the essential information from the input sequence and serves as the basis for generating the output sequence. The decoder component takes the context vector produced by the encoder and generates the output sequence. In ASR, this output sequence consists of characters or words representing the recognized speech. The decoder is typically implemented as a recurrent neural network (RNN), such as a Long Short-Term Memory (LSTM) or Gated Recurrent Unit (GRU) network, or it could be a transformer-based architecture. During training, the model learns to map input sequences to their corresponding output sequences by minimizing a suitable loss function, such as cross-entropy loss. This is typically done using techniques like backpropagation through time (BPTT) or teacher forcing, where the model is trained to predict the next token in the output sequence given the previous tokens. Thereafter, the trained model is used to transcribe unseen speech input. The encoder processes the input sequence to produce the context vector, which is then fed into the decoder to generate the output sequence. In some cases, beam search Szűcs and Huszti (2019); Li et al. (2018) or other decoding strategies may be used to improve the quality of the generated output.
In Chiu et al. (2018), the authors explored various structural and optimization enhancements to their LAS Sequence to Sequence model, resulting in significant performance improvements. They introduce several structural enhancements, including the utilization of word piece models instead of graphemes and the incorporation of a multi-head attention architecture, which outperforms the commonly used single-head attention mechanism. Additionally, they investigate optimization techniques such as synchronous training, scheduled sampling, label smoothing, and minimum word error rate optimization, all of which demonstrate improvements in accuracy. The authors present experimental results utilizing a unidirectional LSTM encoder for streaming recognition. On a 12,500-hour voice search task, they observe a decrease in Word Error Rate (WER) from 9.2% to 5.6% with the proposed changes, while the best-performing conventional system achieves a WER of 6.7%. Moreover, on a dictation task, their model achieves a WER of 4.1%, compared to 5% for the conventional system. Similarly, the work of Prabhavalkar et al. Prabhavalkar et al. (2017) investigated a number of sequence-to-sequence methods in automatic speech recognition. These included the RNN transducer (RNN-T), attention-based models, a new model that augments the RNN-T with attention, and a Connectionist Temporal Classification (CTC) trained system that directly outputs grapheme sequences. According to their research, sequence-to-sequence approaches can compete on dictation test sets against state-of-the-art when trained on a large volume of training data. Even though deep learning has achieved state-of-the-art in speech recognition, an area that still calls for attention is speech-to-speech translation. This is because the present deep learning based speech-to-speech translation systems operate by translating sentences individually, disregarding any contextual information from preceding sentences. While research on contextual understanding has been ongoing for years, challenges persist regarding its practicality and processing efficiency, since translation typically relies on the surrounding words for context.
4.2.3 Electrocardiogram (ECG) Classification
Disorders pertaining to the heart or blood vessels are collectively referred to as Cardiovascular Diseases (CVD) Liu et al. (2021a). According to the American Heart Association’s 2023 statistics, CVD has emerged as the leading cause of death worldwide. In 2020, 19.05 million deaths were recorded from CVD globally, which signifies an increase of 18.71% from 2010, and it is believed that this number will rise to 23.6 million by 2030 Tsao et al. (2023). Blood clots and vascular blockages caused by CVDs can cause myocardial infarction, stroke or even death Liu et al. (2021a). Generally, early diagnosis has been shown to reduce the mortality rate of CVDs, and Electrocardiogram (ECG) signals play a crucial role in diagnosing various cardiac abnormalities and monitoring heart health. However, ECG signal has characteristics of high noise and high complexity, making it time-consuming and labor-intensive to identify certain diseases using traditional methods. The traditional approach is tedious and requires the expertise of a medical specialist. Over the past decades, the task of Long-term ECG recording classification has been significantly facilitated for cardiologists through the adoption of computerized ECG recognition practices. Throughout this period, feature extraction methods have predominantly relied on manual techniques, encompassing diverse approaches such as wave shape functions Llamedo and Martínez (2011), wavelet-based features Mathews et al. (2018), ECG morphology zhu et al. (2019), hermite polynomials Desai et al. (2021), and Karhunen-Loeve expansion of ECG morphology Crippa et al. (2015), among others. These extracted features are subsequently subjected to classification using various machine learning algorithms.
More recently, the advent of deep learning has revolutionized the field by enabling automatic feature learning directly from ECG signals. This advancement holds significant promise in the realm of automated ECG classification, offering clinicians a tool for swift and accurate diagnosis. Based on this, several deep learning architectures have been proposed for feature learning of ECG signals. For instance, Acharya et al. Acharya et al. (2017) developed a 9-layer CNN model to automatically identify five categories of heartbeats in ECG signals. A similar model was also developed in Baloglu et al. (2019), However, ECG signals often vary significantly in length, as they may contain different numbers of heartbeats. CNNs typically require fixed-length inputs, which may necessitate preprocessing steps such as padding or truncation, potentially losing important temporal information. For this reason, several architectures have leveraged RNN in ECG classification, as seen in Singh et al. (2018), Prabhakararao and Dandapat (2020) and Wang et al. (2023b), among others. While RNNs are capable of handling sequential data, they also have limitations in capturing local patterns or short-term dependencies effectively. In ECG signals, local features such as specific waveforms or intervals can be crucial for classification. For this reason, recent works have proposed hybrid models which combine the strengths of both CNNs and RNNs to overcome some of these limitations Sowmya and Jose (2022).
The work of Rai et al. Rai and Chatterjee (2022) developed a hybrid CNN-LSTM network to evaluate the optimum performing model for myocardial infarction detection using ECG signals. The authors then experimented on 123,998 ECG beats obtained from the PTB diagnostic database (PTBDB) and MIT-BIH arrhythmia database (MITDB), and the result showed that by combining the capabilities of both CNN and LSTM, improved classification accuracy can be achieved. Also, in Banerjee et al. (2020), a CNN architecture was developed to extract morphological features from ECG signals. For the purpose of determining the degree of heart rate variability, another composite structure was designed using LSTM and a collection of manually created statistical features. Following that, a hybrid CNN-LSTM architecture is built using the two independent biomarkers to classify cardiovascular artery diseases, and experiments were carried out on two distinct datasets. The first is a partly noisy in-house dataset collected using an inexpensive ECG sensor, and the other is a corpus taken from the MIMIC II waveform dataset. The hybrid model proposed in the work achieved an overall classification accuracy of 88% and 93%, respectively, which surpasses the performance of standalone architectures.
An automated diagnosis method based on Deep CNN and LSTM architecture was presented in Kusuma and Jothi (2022) to identify Congestive Heart Failure (CHF) from ECG signals. Specifically, CNN was used to extract deep features, and LSTM was employed to exploit the extracted features to achieve the CHF detection goal. The model was tested using real-time ECG signal datasets, and the results showed that the AUC was 99.9%, the sensitivity was 99.31%, the specificity was 99.28%, the F-Score was 98.94%, and the accuracy was 99.52%. However, since ECG signals can vary in length due to differences in recording durations or patient conditions. LSTMs are capable of handling variable-length sequences, but traditional CNNs typically require fixed-length inputs. Therefore, fusing these features effectively in a hybrid model can be challenging. Also, Hybrid CNN-RNN models can be computationally intensive, especially when processing long ECG sequences or large datasets. For this reason, recent research works have proposed the use of attention mechanisms to reduce the computational burden by enabling the model to selectively attend to informative features, focusing computational resources where they are most needed. Likewise, attention mechanisms can enable the model to attend to informative segments of the ECG signal, regardless of their length, allowing for more flexible processing of variable-length sequences.
Several researchers have leveraged attention mechanisms in standalone and hybrid architectures for improved performance. For instance, in the work of Chun-Yen et al. Chen et al. (2022a), CNN layers were used to extract main features, while LSTM and attention were included to enhance the model’s feature learning capabilities. Experiments on a 12-lead KMUH ECG dataset showed that the model had high recognition rates in classifying normal and abnormal ECG signals, compared to hybrid models without attention mechanisms. Wang et al. Wang et al. (2021) presented a 33-layer CNN architecture with non-local convolutional block attention module (NCBAM). To extract the spatial and channel information, preprocessed ECG signals were first fed into the CNN architecture. A non-local attention further captured long-range dependencies of representative features along spatial and channel axes. Similarly, a spatio-temporal attention-based convolutional recurrent neural network (STA-CRNN) was presented in Zhang et al. (2020a) with the aim of concentrating on representative features in both the spatial and temporal dimensions. The CNN subnetwork, spatiotemporal attention modules, and RNN subnetwork made up the STA-CRNN and according to findings, the STA-CRNN model was able to classify eight different forms of arrhythmias and normal rhythm with an average F1 score of 0.835.
Combining hybrid deep learning models with attention mechanisms for ECG feature learning is a promising approach that has already shown potential in ECG feature learning, according to reviewed literature. Future research can further explore semi-supervised and self-supervised learning techniques to leverage large amounts of unlabeled ECG data. This could involve pre-training models on large-scale unlabeled datasets using self-supervised learning objectives. Also, deep learning models have been leveraged in the generation of synthetic ECG signals to augment real signals, as seen in Zhu et al. (2019a) where a GAN model was developed to generate ECG signals that correspond with available clinical data. The GAN model used two layers of BiLSTM networks for the generator and CNN for the discriminator, and trained using the 48 ECG recordings of different users from the MIT-BIH dataset. The authors then compared their model with a Recurrent neural network autoencoder (RNN-AE) model and a recurrent neural network variational autoencoder (RNN-VAE) model, and the results showed that their model exhibited the fastest convergence of its loss function to zero. Future research can also incorporate attention mechanisms into hybrid GAN architectures to improve the quality of generated signals. Likewise, real-time detection of heart diseases is paramount, future work can develop efficient algorithms for real-time processing of ECG data. This could involve optimizing existing architectures and leveraging hardware acceleration techniques to enable real-time inference on resource-constrained devices such as wearable sensors and implantable devices.
4.2.4 Electroencephalography (EEG) Classification
Three-dimensional scalp surface electrode readings provide a dynamic time series that is called Electroencephalogram (EEG) signal Schirrmeister et al. (2017). Brain waves obtained from an EEG can effectively depict both the psychological and pathological states of a human. The human brain is acknowledged to be a fascinating and incredibly complicated structure. Numerous brain signals, including functional magnetic resonance imaging (fMRI), near-infrared spectroscopy (NIRS), electroencephalograms (EEGs), and functional near-infrared spectroscopy (fNIR), among others have been collected and used to study the brain Gao et al. (2021b). Due to the EEG’s non-invasive, affordable, accessible, and excellent temporal resolution characteristics, it has become the most utilised approach. However, the signal-to-noise ratio of EEG signal is low, meaning that sources with no task-relevant information frequently have a stronger effect on the EEG signal than those that do. These characteristics often make end-to-end feature learning for EEG data substantially more challenging Schirrmeister et al. (2017). Based on this, several methods have been leveraged for improved feature extraction in EEG signals across several domains including Motor imagery Ang and Guan (2017), anxiety disorder Shen et al. (2022), epileptic seizure detection Boonyakitanont et al. (2020), sleep pattern analysis and disorder detection Sharma et al. (2021); Vaquerizo-Villar et al. (2023), and Alzheimer’s disease detection Modir et al. (2023), and many others.
EEG Motor Imagery (MI) is a technique used to study brain activity associated with the imagination of movement. It involves recording electrical activity generated by the brain through electrodes placed on the scalp. MI tasks typically involve imagining performing a specific motor action, such as moving a hand or foot, without physically executing the movement, and has been leveraged in smart healthcare applications such as post-stroke rehabilitation and mobile assistive robots, among others Altaheri et al. (2023). Prior to the advent of deep learning, motor imagery EEG data are passed through various steps before classification using traditional ML techniques. Pre-processing, feature extraction, and classification are the three primary stages that traditional approaches usually take while processing MI-EEG signals. Pre-processing includes a number of operations, including signal filtering (choosing the most valuable frequency range for MI tasks), channel selection (identifying the most valuable EEG channels for MI tasks), signal normalisation (normalising each EEG channel around the time axis), and artefact removal (removing noise from MI-EEG signals). Independent component analysis (ICA) is the most often utilised technique for removing artefacts Brunner et al. (2007); Delorme et al. (2007); Jafarifarmand and Badamchizadeh (2019). In contrast to the traditional approach, deep learning architectures can automatically extract complex features from raw MI-EEG data without the need for laborious feature extraction and pre-processing. Based on this, several deep learning architectures have been proposed for MI-EEG feature learning, as seen in Zhang et al. (2019b), Kumar et al. (2016) and Tibrewal et al. (2022), among others. For instance Schirrmeister et al. (2017), categorized MI-EEG signals using three CNNs with varying architectures, and the number of convolutional layers varied from two layers to a five-layer deep ConvNet to a thirty-one-layer residual network.
In Dai et al. Dai et al. (2019), the authors proposed an approach for classifying MI-EEG signals which blend variational autoencoder with CNN architecture. The VAE decoder was used to fit the Gaussian distribution of EEG signals, and the time, frequency, and channel information from the EEG signal were combined to create a novel representation of input, and the proposed CNN-VAE method was optimised for the input. Experiments showed that by combining both deep learning architectures, improved features were learnt, which led to a high classification performance on the BCI Competition IV dataset 2b. Li et al. Li et al. (2017) employed optimal wavelet packet transform (OWPT) for the generation of feature vectors from MI-EEG signals. These vectors were then utilized to train an LSTM network which demonstrated satisfactory performance on dataset III from the BCI Competition 2003. However, the model has excessively intricate structure. To address this, Feng et al. Li et al. (2020a) introduced a technique that merges continuous wavelet transform (CWT) with a simplified convolutional neural network to enhance the accuracy of recognizing MI-EEG signals. By employing CWT, MI-EEG signals were converted into time-frequency image representations. Subsequently, these image representations were fed into the SCNN for feature extraction and classification. Experiments on the BCI Competition IV Dataset 2b demonstrate that, on average, the classification accuracy across nine subjects reached 83.2%. However, the computational complexity of the model was quite high, due to the processing of time-frequency image representations. The conversion of MI-EEG signals into time-frequency images using CWT requires significant computational resources.
The authors in Hwang et al. (2023) introduced a classification framework based on Long Short-Term Memory (LSTM) to improve the accuracy of classifying four-class motor imagery signals from EEG. The authors sliding window technique to capture time-varying EEG signal data, and employed an overlapping-band-based Filter Bank Common Spatial Patterns (FBCSP) method to extract subject-specific spatial features. Experiments on the BCI Competition IV dataset 2a, showed that their model achieved an average accuracy of 97%, compared to existing methods. Also, in the classification of Alzheimer’s disease, Zhao et al. Zhao and He (2015) employed a deep learning network to analyse EEG data. The deep learning model was evaluated on a dataset that consist of fifteen (15) patients with clinically confirmed Alzheimer’s disease and fifteen (15) healthy individuals, and results showed that improved features were learnt and compared the results to the traditional methods. This has prompted the use of deep learning in Alzheimer’s disease detection, as seen in Xia et al. (2023), where the authors used CNN for diagnosing Alzheimer’s Disease. To address challenges posed by limited data and overfitting in deep learning models designed for Alzheimer’s Disease detection, the authors explored the use of overlapping sliding windows to augment the EEG data collected from 100 subjects (comprising 49 AD patients, 37 mild cognitive impairment patients, and 14 healthy controls subjects). After assembling the augmented dataset, a modified Deep Pyramid Convolutional Neural Network (DPCNN) was used to classify the enhanced EEG signals. In epilepsy detection, Hermawan et al. (2024) developed three deep learning architectures (CNN, LSTM, and hybrid CNN-LSTM), with each model chosen for its effectiveness in handling the intricate characteristics of EEG data. Each architecture offers distinct advantages, with CNN excelling in spatial feature extraction, LSTM in capturing temporal dynamics, and the hybrid model combining these strengths. The CNN model, consisting of 31 layers, attained the highest accuracy, achieving 91% on the first benchmark dataset and 82% on the second dataset using a 30-second threshold, selected for its clinical significance.
In the work of Abdulwahhab et al. Abdulwahhab et al. (2024), EEG waves’ time-frequency image and raw EEG waves served as input elements for CNN and LSTM models. Two signal processing methods, namely Short-Time Fourier Transform (STFT) and CWT, were employed to generate spectrogram and scalogram images, sized at 77 × 75 and 32 × 32, respectively. The experimental findings demonstrated detection accuracies of 99.57% and 99.26% for CNN inputs using CWT Scalograms on the Bonn University dataset and 99.57% and 97.12% using STFT spectrograms on the CHB-MIT dataset. Similarly, in emotion recognition, several deep learning models have been leveraged with EEG signals. For instance, in Pandey and Seeja (2022), a subject-independent emotion recognition model was proposed, which utilizes Variational Mode Decomposition (VMD) for feature extraction and DNN as the classifier. Evaluation against the benchmark DEAP dataset demonstrates superior performance of this approach compared to other techniques in subject-independent emotion recognition from EEG signals. Also, some researchers have also combined EEG signals with facial expression and speech in emotion recognition, as seen in Hassouneh et al. (2020), Pan et al. (2023), and Wang et al. (2023c), among others. However, EEG signals can vary significantly across individuals, making it challenging to generalize models across different subjects. Future models could explore methods for adapting or personalizing models to account for inter-subject variability and improve performance on individual subjects. Also, EEG electrodes cover only a fraction of the brain’s surface, resulting in limited coverage of neural activity. Deep learning models could investigate strategies to infer activity from unobserved brain regions or integrate information from multiple modalities to provide more comprehensive coverage. These areas can still be further explored.
4.2.5 Finance
Over the past few decades, computational intelligence in finance has been a hot issue in both academia and the financial sector Ozbayoglu et al. (2020). Deep learning, especially RNN models have gained significant traction in the field of finance due to its ability to handle sequential data, since financial data often exhibit sequential dependencies, such as time series data for stock prices or historical transaction data. Within the financial industry, researchers have developed deep learning models for stock market forecasting Singh and Srivastava (2017), algorithmic trading Lei et al. (2020), credit risk assessment Shen et al. (2021), portfolio allocation Wang et al. (2020b), asset pricing Chen et al. (2024), and derivatives markets Ahnouch et al. (2023), among others and these models are intended to offer real-time operational solutions. In exchange rate prediction, Sun et al. Sun et al. (2020) developed an ensemble deep learning technique known as LSTM-B by combining a bagging ensemble learning algorithm with a long-short term memory (LSTM) neural network to increase the profitability of exchange rate trading and produce accurate exchange rate forecasting results. In comparison to previous methodologies, the authors’ estimates proved to be more accurate when they looked at the potential financial profitability of exchange rates between the US dollar (USD) and four other major currencies: GBP, JPY, EUR, and CNY.
The authors in Abedin et al. (2021) proposed a Bi-LSTM-BR technique, which combined Bagging Ridge (BR) regression with Bi-LSTM as base regressors. The pre-COVID-19 and COVID-19 exchange rates of 21 currencies against the USD were predicted using the Bi-LSTM BR, and experiments showed that the proposed method outperformed ML algorithms such as DT and SVM. However, exchange rate data can be noisy and subject to non-stationarity, which can pose challenges for predictive modelling. While bagging techniques can help mitigate the effects of noise to some extent, they may struggle to capture long-term trends or sudden shifts in the data distribution, leading to suboptimal performance. To address this, Wang et al. Wang et al. (2023a) presented an approach for one-day ahead of time exchange rate prediction that concurrently considers both supervised and unsupervised deep representation features to enhance Random Subspace. Two crucial phases in the SUDF-RS technique are feature extraction and model building. First, LSTM and deep belief networks, respectively, extract the supervised and unsupervised deep representation features. To produce high-quality feature subsets, an enhanced random subspace approach was created that integrates a random forest-based feature weighting mechanism. Then, the matching base learner is trained using each feature subset, and the final outcomes are generated by averaging the outcomes of each base learner. Experiments on EUR/USD, GBP/USD and USD/JPY showed that improved accuracy was achieved using the model.
In stock market prediction, several deep learning architectures have been proposed in the literature. For instance, Nikou et al. (2019), conducted a comparative study between the ANN, SVR, RF and an LSTM model. As compared to the other models discussed in the study, the LSTM model outperformed the others in predicting the closing prices of iShares MSCI United Kingdom. Similarly, using stock market historical data and financial news, Cai et al. Cai et al. (2018) used CNN and LSTM forecasting methods to generate seven prediction models. The seven models were then combined into a single ensemble model in accordance with the ensemble learning approach to create an aggregated model. However, the accuracy of all the models’ predictions was low. Gudelek et al. Gudelek et al. (2017) proposed a CNN model which used a sliding window technique and created pictures by capturing daily snapshots within the window’s bounds. With 72% accuracy, the model was able to forecast the prices for the following day and was able to generate 5 times the starting capital. In Eapen et al. Eapen et al. (2019), a CNN and Bi-LSTM model with numerous pipelines was proposed, utilising an SVM regressor model on the S&P 500 Grand Challenge dataset, and results showed enhanced prediction performance by over a factor of 6% compared to baseline models. As presented, deep learning has undeniably achieved state-of-the-art performance across various domains within finance. However, due to the sensitive nature of financial research, future work can focus on enhancing the interpretability of deep learning models in financial predictions. Researchers should explore techniques to explain the predictions of models, to improve trust and understanding of model decisions, which is essential for adoption in finance.
4.3 Natural Language Processing
Natural language processing (NLP) refers to the field of artificial intelligence that concerns with enabling computers to process, analyze and interpret human languages to extract useful information. Some of the common tasks in NLP are machine translation, text classification and text generation. Deep learning has been widely applied to solve real-world NLP problems. This section presents the recent advancements in deep learning models that have been designed for NLP over the past few years.
4.3.1 Text Classification
Text classification known as text categorization, is a task that involves assigning predefined categories or labels to a piece of text based on its content. The task is commonly used in various applications such as document classification, sentiment analysis and spam filtering. Numerous deep learning models have been proposed for text classification in the past few decades, and multilayer perceptron is one of the earliest architectures adopted to classify documents Calvo and Ceccatto (2000); Yu et al. (2008). The model typically has a single hidden layer with a number of units between 15 and 150. Text data is inherently sequential, as it is composed of a series of words and symbols arranged in a specific order. This property makes RNN and its variants particularly well-suited for processing and analyzing text data. In Arevian (2007), RNN with two hidden layers, each with 6 units is used to classify news documents into eight classes. A study was conducted to investigate the variants of RNN i.e. LSTM and GRU for text classification Huang and Feng (2020). The input to the model is a sequence of words of fixed length. The input sequence is also sliced into smaller subsequences of fixed length and passed to an independent model for parallelization. A convolutional layer can extract local features, allowing the model to leverage hierarchical temporal information in textual data. A hybrid model of convolutional and LSTM architecture is proposed for text classification Wang et al. (2019). Two parallel convolutional layers are used to extract features from word embeddings, followed by max-pooling layers to reduce the feature dimensions. The reduced features are then concatenated and passed to LSTM for prediction.
Although CNN and RNN provide excellent results on text classification tasks, the models lack the ability to attend to specific words based on their importance and context. To address this limitation, attention mechanism is incorporated into the model to focus on the important features, enhancing the text classification accuracy. In Liu and Guo (2019), two attention modules are introduced to capture the contextual information of the feature sequence extracted by bi-directional LSTM. The first attention module attends the sequence in forward direction while the backward sequence is attended by the second attention module. The convolutional layers are used before the bi-directional LSTM to extract features from the word embedding. The attention modules require sequential processing using RNN-based architecture such as LSTM and GRU which may lead to information loss and distorted representations, particularly in long sequence. Furthermore, the attention modules focus on inter-sequence relationships between the input sequence and the target, ignoring the intra-sequence relationships or the dependencies between the words. In Lin et al. (2017b), the deep learning model is integrated with self-attention to capture the intra-sequence relationships between the features in the sequence. A multilayer of bi-directional LSTMs is utilized to extract feature sequence from the word embedding before the self-attention module attends the feature sequence to compute the attention weights. To further improve the overall performance, a multichannel features consisting of three input pipelines is introduced Li et al. (2020b). Each pipeline concatenates the word vector with a feature vector derived from the input sequence such as the word position, part-of-speech and word dependency parsing. The input pipeline is connected to bi-directional LSTM, followed by a self-attention module to learn the dependencies between the features in the sequence.
The transformer is a deep learning architecture that transforms sequential data using self-attention mechanisms, allowing long-range dependencies and complex patterns to be captured. The architecture is the basis of various advanced deep learning models and the Bi-directional Encoder Representations from Transformers popularly known as BERT is one of the examples that leverage transformer for pre-training on large scale textual data Devlin et al. (2018). BERT is a bi-directional transformer encoder that is designed for various NLP tasks, capable of capturing the contextual information from both preceding and succeeding words in the input sequence. Several improvements have been made to BERT to enhance its overall performance such as ALBERT Lan et al. (2019), RoBERTa Liu et al. (2019b) and DeBERTa He et al. (2020a). The improvements are centered around refining the pre-training approaches such as dynamic masking of the training instances, training with a block of sentences and representing each input word using two vectors, both content and position of the word. Most of the recent works leverage BERT and its variants to capture effective feature representation of the input sequence. In Rodrawangpai and Daungjaiboon (2022), BERT and its variants are leveraged to capture the long-range dependencies of the input tokens. The features are then passed to a layer normalization and a linear fully-connected layer with dropout for classification. Similar work is reported in Murfi et al. (2024) whereby BERT is used to extract the features and the features are then passed to a hybrid of convolutional and recurrent neural networks. The traditional machine learning algorithms have been used to classify the features extracted by BERT Hao et al. (2023). The study shows machine learning algorithms can effectively leverage the rich contextual features extracted by BERT for downstream classification tasks.
In text classification, the text labels can help in capturing the words relevant to the classification. The label-embedding attentive model is one of the earliest attempts to joint learn the label and word embeddings in the same latent space and measure the compatibility between labels and words using cosine similarity Wang et al. (2018a). The joint embedding allows the model to capture more effective text representations, increasing the overall performance of the model. LANTRN is a deep learning model that leverages label embedding extracted by BERT and entity information e.g. person name and organization name for text classification Yan et al. (2023). The entity recognition module is based on bi-directional LSTM and conditional random field layers to calculate the probability of each word in each entity label. The model introduces a label embedding bi-directional attention to learn the attention weights of token-label and sequence-label pairs. Furthermore, a transformer is introduced to learn local short-term dependencies of multiple short text sequences and long-term dependencies of the input sequence. Aspect refers to a specific attribute of an entity within the text and incorporating this information enhances the model’s understanding of the nuances of the text. BERT-MSL is a multi-semantic deep learning model with aspect-aware enhancement and four input pipelines: left sequence, right sequence, global sequence and aspect target Zhu et al. (2023). The aspect-aware enhancement module takes the features extracted by BERT, and performs average pooling followed by a linear transform. Then the output is concatenated with the outputs produced by the local and global semantic learning modules. The concatenated features are then jointly attended by a multi-head attention for text classification. Figure 24 shows the architecture of BERT-MSL.
Figure 24: The architecture of BERT-MSL Zhu et al. (2023).
4.3.2 Neural Machine Translation
Neural machine translation (NMT) refers to the automated process of translating text from one language to another language. Numerous deep learning models have been proposed for NMT which can be categorized into RNN-based and CNN-based models. One of the first successful RNN-based models is the encoder-decoder Cho et al. (2014); Sutskever et al. (2014). The model consists of two connected subnetworks (the encoder and the decoder) for modelling the translation process as shown in Figure 25. The encoder reads the source sentence word by word and produces a fixed-length context vector (final hidden state). This process is known as source sentence encoding as shown in the figure. Given the context vector, the decoder generates the target sentence (translation) word by word. This modelling of the translation can be seen as a mapping between the source sentence to the target sentence via the intermediate context vector in the semantic space. The context vector represents the summary of the input sequence’s semantic meaning, providing a compressed representation that captures the essence of the source sentence. However, the compression process can sometimes result in the loss of information especially those early in the sequence. Bi-directional RNN may mitigate the loss of information by modelling the sequence in reverse order. However, the problem can still persist, particularly in cases where the input is a long sequence.
Figure 25: The architecture of an encoder-decoder Stahlberg (2020).
Attention mechanism was introduced to solve the problem of learning long input sequences Bahdanau et al. (2014). Attention alleviates this issue by attending on different words of the input sequences when predicting the target sequences at each time step. Unlike the standard encoder-decoder model, attention derives the context vector from the hidden states of both the encoder and decoder, and the alignment between the source and target. This mechanism allows the model to focus on the important words, increasing the overall accuracy of the translation. Several alignment score functions have been proposed for calculating the attention weights. Some of the popular functions are additive Bahdanau et al. (2014), dot-product, location-based Luong et al. (2015), and scaled dot-product Vaswani et al. (2023). The attention weights are calculated by attending to the entire hidden states of the encoder. This attention, also known as global attention, is computationally expensive. Instead of attending to all hidden states, local attention attends to a subset of hidden states, thus reducing the computational cost Luong et al. (2015). Google Neural Machine Translation is a popular encoder-decoder model with an attention mechanism that significantly improves the accuracy of machine translation Wu et al. (2016). As shown in Figure 26, the model consists of a multilayer of LSTMs with eight encoder and decoder layers and an attention connection between the bottom layer of the decoder to the top layer of the encoder. Furthermore, to deal with the challenging words to predict, a word is tokenized into subwords e.g. feud is broken down into fe and ud, allowing the model to generalize well to new and uncommon words. A year later, the self-attention mechanism was proposed, significantly improving the overall accuracy of machine translation Vaswani et al. (2023). Self-attention, also known as intra-attention, allows the deep learning model to capture the dependencies between the input words. The self-attention mechanism is the fundamental building block of the transformer model, which has since become a cornerstone in natural language processing and other domains.
Figure 26: The architecture of Google Neural Machine Translation Wu et al. (2016).
Despite the success of transformer, the model falls short in capturing nuances of human language and struggles with tasks requiring deeper understanding of context. This can be especially challenging when the tasks involve formality, colloquialism, and subtle cultural references that may not directly equivalent in the target language, resulting in inaccurate translation or losing the original meaning. One of the approaches to include context into the input sequence is concatenating the current source sentence with the previous (context) sentences and feeding the whole input to the transformer Lupo et al. (2023). The model is trained to predict the translated sentence including the context translation. At inference time, only the translation is considered while the context translation is discarded. Furthermore, the approach encodes the sentence position and segment-shifted position to improve the distinction between current sentences and context sentences. In Rippeth et al. (2023), the source sentence is prefixed with the summary of the document to contextualize the input sentence. The summary is the set of salient words that represents the essence of the document, resolving ambiguity associated with the translation. A study was conducted to determine the optimal technique of aggregating contextual features Wu et al. (2022). Three techniques were studied namely concatenation mode, flat mode and hierarchical mode, and the experimental results show that concatenation mode achieved the best results. In Kim et al. (2023), a training method is introduced to train the deep learning machine translation model to generate translation involving honorific words. The training method indicates the honorific context in the target sentence using an honorific classifier to guide the model to attend to the related tokens. Unlike other studies where the context features are included by concatenation, the training method assigns weights to the context tokens indicated by the honorific classifier. This allows the model to generate a more accurate translation with honorifics. Finally, the performance of transformer relies on large-scale training data. However, for the vast majority of languages, only limited amounts of training data exist. To mitigate this problem, recent studies introduce shallow transformer architectures Gezmu and Nürnberger (2022), explore the effect of hyperparameter finetuning Araabi and Monz (2020), exploiting monolingual corpus to enhance the bilingual dataset for model training Li et al. (2024) and leveraging visual input as contextual information for the translation task Meetei et al. (2023).
4.3.3 Text Generation
Text generation refers to the process of creating texts based on a given input whereby the input can be in the form of texts, images, graphs, tables or even tabular data. Due to the various forms of inputs, text generation has a wide range of applications, including creative writing, image captioning and music generation. This section focuses on the progress made in text-to-text generation tasks such as question answering, dialogue generation and text summarization. The recurrent neural network and its variants play an important role in text generation tasks for their strong ability to model sequential data. One of the earliest works on question answering is based on the RNN-based encoder-decoder model whereby the encoder takes the question embedding and processes it using bi-directional LSTM, and the decoder generates the corresponding answer Nie et al. (2017). Additionally, to prevent semantic loss and enable the model to focus on the important words in the input sequence, a convolution operation is applied to the word embedding, and an attention mechanism is then used to attend to the output of the convolution operation. Similar work is reported in Yin et al. (2015) in which a knowledge-based module is introduced to calculate the relevance score between the question and the relevant facts in the knowledge base. This improves the text (answer) generation by the decoder. Another work is described in Li et al. (2016) where an encoder-decoder with attention for dialogue generation is optimized using reinforcement learning. The model is first trained in supervised learning manner and then improved using the policy gradient method to diversify the responses. Ambiguous content in question answering sentences is a challenge in text generation and can lead to incorrect and uncertain responses. Cross-sentence context aware bi-directional model introduces a parallel attention module to compute the co-attention weights at the sentence level, accounting for the relationships and similarities in the question and the answer Wu et al. (2020).
The transformer has been leveraged for text generation tasks. An incremental transformer-based encoder is proposed to incrementally encode the historical sequence of conversations Li et al. (2019b). The decoder is a two-pass decoder that is based on the deliberation network, generates the next sentence. The first pass focuses on contextual coherence of the conversations while the second pass refines the output of the first pass. BERT and ALBERT have been used as pre-trained models for question answering task Alrowili and Vijay-Shanker (2021). The study found that the performance of the models is sensitive to random assignment of the initial weights especially on small datasets Alrowili and Vijay-Shanker (2022). T-BERTSum is a model based on BERT, designed to address the challenge of long text dependence and leveraging latent topic mapping in text summarization Ma et al. (2021). The mode integrates a neural topic module to infer topics and guide summarization, uses a transformer network to capture long-range dependencies and incorporates multilayers of LSTM for information filtering. Exploiting domain knowledge is essential in reducing the semantic gap between the deep learning models and the text corpus. KeBioSum is a knowledge infusion framework to inject domain knowledge into the pre-trained BERTs for text summarization Xie et al. (2022). In the framework, the relevant information is detected and extracted from the domain knowledge, generating label sequences of the sentences. The label data is then used to train the text summarization model using discriminative and generative training approaches, infusing the knowledge into the model.
5 Challenges and Future Directions
5.0.1 Availability and Quality
Building and employing deep learning models face several challenges. The training of deep learning requires a large number of instances (examples) to achieve high accuracy and generalization Munappy et al. (2022). Furthermore, the complexity of deep neural networks may lead to overfitting, where the model performs well on training data but fails to generalize on new, unseen data. This phenomenon frequently arises when the models is trained on insufficient data, highlighting the importance of diverse and extensive datasets. However, the data collection and annotation are time consuming and often require domain experts, specialized training and standardization Luca et al. (2022). Moreover, this process is prone to error and has the risk of introducing biases into the dataset which can significantly impact the performance of the trained model. One of the approaches to address this issue is transfer learning. Transfer learning involves the use of a deep learning model (known as pre-trained model) that is trained on a large dataset for solving a specific task (with a small dataset) Zhuang et al. (2020). The pre-trained model serves as a basis for the model training by fine-tuning the weights of the pre-trained model and adapting it to the new prediction task. This approach helps to mitigate the lack of training data in the target domain. Furthermore, transfer learning reduces computational resources required to train the model and helps faster convergence.
Another approach that can be employed to address the lack of data is data augmentation. Data augmentation is a convenient method that increases the number of instances by performing transformation functions on the existing instances without changing the labels Mumuni and Mumuni (2022). In the domain of computer vision, image transformation such as rotation, translation and cropping. However, it is important to consider the output of the transformation because the resultant may not represent the actual data. For example, flipping or adding noise to a signal may introduce distortion or changing the characteristics (trend, seasonality and cyclic variations) of the signal. Thus, careful consideration must be given to ensure that the generated instances still accurately represent the underlying patterns present in the data. Data augmentation can also be realized by generating synthetic data to supplement the training set. Synthetic data is artificially created data that resembles real data but is generated using statistical methods or deep generative models Hu et al. (2023); Murtaza et al. (2023). The generated data can complement the less-diverse, limited datasets, providing a broader range of examples for the model to learn from. However, generating synthetic data that accurately reflects the characteristics of the real-world data is challenging. Careful consideration must be given to the choice of models and parameters used to ensure the synthetic data is realistic and representative of the real-world data.
5.0.2 Interpretability and Explainability
Interpretability and explainability is crucial for building trust and understanding how predictive models make decisions especially in high-stake applications such as healthcare and medical image analysis Tonekaboni et al. (2019). However, as deep learning models become more intricate and complex with numerous layers, subnetworks and a large number of parameters, the models are often perceived as a “black box” and difficult to explain in terms of decision-making processes. Therefore, it is crucial for the researchers to focus on methods that provide insights into how a deep learning model performs the prediction and how its decisions are influenced by the input data, making it more transparent and trustworthy. Numerous methods have been proposed for interpreting and explaining the decisions of deep learning models which can be categorized into visualization (feature attribution), model distillation and intrinsic (explainable by itself). Visualization methods involve the use of scientific visualization such as saliency maps or heatmaps to express the explanation by highlighting the degree of association between the inputs and the predictions Tjoa et al. (2023). The heatmaps identify the saliency of the input features influencing the model’s predictions. The visualization approach is simple and intuitive and can be applied to tabular data and image data. Furthermore, it can be used to identify and debug issues in deep learning models, leading to improved performance and robustness.
Model distillation is an approach to approximating a complex model by fitting a simpler model using the training set. The simpler model is built typically using a simpler or interpretable algorithm such as linear regression, decision tree or rule-based methods Li and Shen (2024). In this approach, the simpler model is trained to resemble the predictive behavior of the complex model. Then, the simpler model may serve as the proxy or surrogate model for explaining the complex model. Model distillation can be used together with visualization to further enhance the interpretability of the complex model Termritthikun et al. (2023). Model distillation seeks explanations of the models that were never designed to be explainable. Ideally, the explanation of a deep learning model’s prediction should be included as part of the model output, or the explanation can be derived from the architecture of the model. This is because an intrinsic model can learn not only the mapping between the input and output, but also generate an explanation of the prediction that is faithful to the model’s behavior. Attention mechanisms are the key to this approach, providing a form of attention weights that can be used to explain why the model made a particular decision Xiong et al. (2022). Another type of intrinsic approach is to train the model to simultaneously perform the prediction task and generate the explanation for its predictions Fernandes et al. (2023). This “additional task” can be in the form of a text explanation or model prototype which embeds the semantic meaning of the prediction. However, the intrinsic approach is more difficult to apply because the user needs additional knowledge and understanding of the model’s architecture and inner workings.
5.0.3 Ethics and Fairness
Deep learning models are increasingly being deployed in making high-stake decision including recruitment Freire and de Castro (2021), criminal justice Dass et al. (2023) and credit scoring Gicić et al. (2023). There are several advantages of deep learning-based systems in which, unlike humans, machines are able to process vast amounts of data and applications quickly and consistently. However, deep learning-based systems have the risk of being prone to biases present in the data used for training which can lead to unfairness and injustice. Numerous efforts have been made to mitigate this issue which can be categorized into modelling bias detection and modelling bias mitigation. Detection of modelling bias refers to the process of identifying and quantifying biases that may present in predictive models. This approach involves the use of statistical analysis, fairness metrics, counterfactual testing and human review to detect bias in the models. For instance, visualization-based methods such as attribution maps are used to indicate which regions are significant to the predictions Schaaf et al. (2021). This in turn can be used to detect and quantify bias using metrics such as Relevance Mass Accuracy, Relevance Rank Accuracy Accuracy and or Area over the perturbation curve (AOPC). In Giloni et al. (2022), two modules are presented for estimating bias in predictive models. The first module utilizes an unsupervised deep neural network with a custom loss function to generate hidden representation of the input data called bias vectors, revealing the underlying bias of each feature. The second module combines these bias vectors into a single vector representing the bias estimation of each feature, achieved by aggregating them using the absolute averaging operation. Bias mitigation refers to the process of reducing the presence of bias in predictive models, which can be done in three stages. The first stage combats bias by modifying the training data, either relabeling the labels or perturbing the feature values Iosifidis et al. (2019); Kehrenberg et al. (2020). The second stage addresses bias during the training of the model by applying regularization terms to the loss function to penalize discrimination. In Jain et al. (2023), a loss function based on bias parity score (BPS) is introduced to measure the degree of similarity of a statistical measure such as accuracy across different subgroups. The BPS term is added to the loss function as a regularizer to the original prediction task. The last stage mitigates bias after the predictive models have been successfully trained. This stage applies post-processing approaches such as reinforcement learning to obtain a fairer model Yang et al. (2023). For instance, the detection of minority classes is rewarded to prevent bias towards the majority class. This allows the model the generalize well across different patient demographics.
5.0.4 Lightweight Deep Learning Models
Even though deep learning architectures have achieved state-of-the-art across various computer vision tasks, they often come with large model parameters Ige and Mohd Noor (2023). The architecture and complexity of a deep learning network determine the number of model parameters. The deeper the network, the larger the number of model parameters. However, deep learning models with large parameters often suffer limitations when deploying on end devices. For instance, a deep learning model developed for security monitoring by analyzing video data using 3D-CNNs might suffer deployment issues when deploying such models on low-resourced systems like smartphones or small-scale IoT devices. Model training and inference for deep learning models with large parameters demands substantial processing power. As the number of parameters increases, so does the computational complexity, resulting in longer model training duration and more hardware needs. Also, large parameter sizes translate to increased memory requirements, limiting their deployment on end devices. This is because these end devices often have battery, processor or memory capacity limitations. To address these challenges, it is important to develop sophisticated but lightweight architectures that can achieve state-of-the-art with few model parameters. Such lightweight models will be characterized by their ability to deliver competitive performance while mitigating computational complexity and memory requirements, making them well-suited for deployment on resource-constrained devices. An approach would be to develop novel lightweight plug-and-play modules that can be plugged to few layered deep learning architectures to improve feature learning without incurring additional model complexity. Other approaches could involve leveraging model compression techniques to reduce the size and computational complexity of deep learning models. Researchers can focus on improving pruning methods Li et al. (2019a), which can identify and eliminate redundant parameters or connections, thereby reducing the model’s footprint without compromising performance. Also, quantization techniques Yang et al. (2019) can be further explored to reduce the precision of weights and activations, therefore, enabling efficient representation with lower memory requirements. Also, knowledge distillation techniques Stanton et al. (2021) can be further investigated to facilitate the transfer of knowledge from a complex teacher model to a simpler student model, therefore, enabling compact yet effective representations. These areas are still open to contributions.
5.0.5 Adversarial Attack and Defense
Adversarial attacks and defense mechanisms in deep learning represent a critical area of research and development, particularly as deep learning models become increasingly integrated into various applications. Adversarial attacks involves the deliberate manipulation of input data to mislead or deceive deep learning models, leading to incorrect predictions or behavior Akhtar and Mian (2018). Szegedy et al. Szegedy et al. (2013) was the first to identify this intriguing shortcoming of deep neural networks in image classification. They showed that even with their great accuracy, deep learning models are surprisingly vulnerable to adversarial attacks that take the form of tiny image changes that are (almost) invisible to human vision systems. A neural network classifier may radically alter its prediction about an image as a result of such an attack. Also, such a model can indicate high confidence in wrong predictions, which can be catastrophic for deep learning models deployed in medical or security fields, among many others. In generative models, several studies have investigated how adversarial attacks affect autoencoders and GANs, as seen in Tabacof et al. Tabacof et al. (2016) where a method to manipulate input images in a way that deceives variational autoencoders into reconstructing a totally different image was introduced. In recent times, the focus of adversarial attack research has been on images, but studies has shown that adversarial attacks are not limited to image data; they can also affect other types of data such as text, signals, audio, and video Zhang et al. (2020b); Jiang et al. (2019); Esmaeilpour et al. (2019). Future research can focus on exploring adversarial attacks in these domains and developing tailored defense mechanisms. Also, researchers can further investigate the practical implications of adversarial attacks in real-world scenarios, such as in autonomous vehicles, medical imaging, and cybersecurity. Understanding the potential impact of adversarial attacks in these applications can inform the development of more robust and secure systems.
6 Conclusions
Deep learning has become the prominent data-driven approach in various state-of-the-art applications. Its importance lies in its ability to revolutionize many aspects of research and industries and tackle complex problems which were once impossible to overcome. Numerous surveys have been published on deep learning, reviewing the concepts, model architectures and applications. However, the studies do not discuss the emerging trends in the state-of-the-art applications of deep learning and emphasise the important traits and elements in the models. This paper presents a structured and comprehensive survey of deep learning, focusing on the latest trends and advancements in state-of-the-art applications such as computer vision, natural language processing, time series analysis and pervasive computing. The survey explores key elements and traits in modern deep learning models, highlighting their significance in addressing complex challenges across diverse domains. Furthermore, this paper presents a comprehensive review of the deep learning fundamentals, which is essential for understanding the core principles behind modern deep learning models. The survey finishes by discussing the critical challenges and future directions in deep learning.
Acknowledgement
This work has been supported in part by the Ministry of Higher Education Malaysia for Fundamental Research Grant Scheme with Project Code: FRGS/1/2023/ICT02/USM/02/2.
Declaration of competing interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Data availability
No data was used for the research described in the article.
[…]